### v3

# On the modified $J$-equation

Abstract.
In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16]. We show that the solvability of the modified $J$-equation is equivalent to the coercivity of the modified $J$-functional on compact Kähler manifolds. For smooth projective toric varieties we establish a Nakai–Moishezon type criterion for the existence of solutions, extending the results of Collins–Székelyhidi [CS17]. As a possible application, by combining our result with Delcroix–Jubert [DJ25], we further discuss a numerical sufficient condition for the existence of extremal Kähler metrics on smooth projective toric varieties.
Key words and phrases:
$J$-equation, coercivity, Nakai–Moishezon criterion
2020 Mathematics Subject Classification:
Primary 53C55; Secondary 35A01
1. Introduction
Let $X$ be an $n$-dimensional compact complex manifold and $\operatorname{Aut}_{\rm red}(X)$ be its reduced automorphism group, i.e. the connected subgroup of $\operatorname{Aut}(X)$ whose Lie algebra consists of holomorphic vector fields which have a zero on $X$. Let $T\subset\operatorname{Aut}_{\rm red}(X)$ a real torus with $\mathfrak{t}:=\operatorname{Lie}(T)$. Let $\omega$, $\widehat{\chi}$ be $T$-invariant Kähler forms, and denote their Kähler classes by $\alpha$ and $\beta$ respectively. Let $\mathcal{H}(X,\widehat{\chi})^{T}$ be the space of $T$-invariant $\widehat{\chi}-$Kähler potentials
$\mathcal{H}(X,\widehat{\chi})^{T}:=\big\{\phi\in C^{\infty}(X;\mathbb{R})^{T}\big|\chi_{\phi}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi>0\big\}.$
[…]
$i_{v}\chi_{\phi}=\frac{\sqrt{-1}}{2\pi}\bar{\partial}\theta_{v}^{X}(\chi_{\phi}),\quad\int_{X}\theta_{v}^{X}(\chi_{\phi})\chi_{\phi}^{n}=0.$ (1.1)
Finding canonical metrics has been one of the central themes in Kähler geometry. Especially several equations involving Hamiltonians have been studied extensively in the last few decades (e.g. [BJ25, FO19, Lah19, Li21, Mab03, TZ00, Zhu00]). Hamiltonians of holomorphic vector fields encode the symmetries of the manifold with respect to a given Kähler metric. When studying canonical metrics, such as Kähler–Einstein or constant scalar curvature Kähler (cscK) metrics, it is natural to take these Hamiltonians into account, since the existence and uniqueness of such metrics are often obstructed by holomorphic automorphisms. By introducing Hamiltonians of holomorphic vector fields, one can treat the variational problem even without a canonical metric.
In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16]:
$\omega\wedge\chi_{\phi}^{n-1}=\frac{1}{n}(c_{X}+\theta_{v}^{X}(\chi_{\phi}))\chi_{\phi}^{n},$ (1.2)
where the constant $c_{X}$ only depends on $\alpha$, $\beta$, and determined by
$c_{X}:=\frac{n\alpha\cdot\beta^{n-1}}{\beta^{n}}.$ (1.3)
Also we note that the constant
$m_{X}:=c_{X}+\min_{X}\theta_{v}^{X}(\chi_{\phi})$ (1.4)
is independent of the choice of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, and the existence of a solution to (1.2) easily leads to $m_{X}>0$ (see Section 2.2). In particular, if $T$ is trivial then the equation (1.2) matches up with the $J$-equation introduced by Donaldson [Don03] and Chen [Che00]. Although the $J$-equation itself is an intriguing equation, it has sometimes been studied in relation to cscK metrics, as it provides the starting point of the continuity method for solving the cscK equation [CC18, CC21a, CC21b]. In the modified case, Li–Shi [LS16] showed that such a solution $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ to (1.2) exists if and only if there exists $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$ such that
$\big(c_{X}+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0,$ (1.5)
where the positivity of (n1,n1)(n-1,n-1)-forms is understood in the sense of [Dem12, Chapter III, Section 1.A]. The solutions to the equation (1.2) are characterized as the critical points of a strictly convex function $J_{v}^{\omega}$ on $\mathcal{H}(X,\widehat{\chi})^{T}$, called the modified $J$-functional. On the other hand, the modified $K$-energy $K_{v_{\rm ext}}$ is a convex functional on the space of Kähler potentials whose critical points are extremal Kähler metrics. Motivated by [LS16], Delcroix–Jubert [DJ25] recently considered a modified version of the Chen–Tian formula, relating these functionals, and proved that the existence of a solution to the modified $J$-equation implies the $T^{\mathbb{C}}$-coercivity of the modified $K$-energy.
For later arguments, it is convenient to extend (1.2) to a more general equation of the form:
$\operatorname{Tr}_{\chi_{\phi}}\omega+b\frac{\omega^{n}}{\chi_{\phi}^{n}}=c+\theta_{v}^{X}(\chi_{\phi}),$ (1.6)
where the constants $b\in\mathbb{R}$, $c>0$ are related with each other by
$b=\frac{c\beta^{n}-n\alpha\cdot\beta^{n-1}}{\alpha^{n}}.$ (1.7)
In most cases, we assume that $b\geqslant 0$, but sometimes allow $b$ to be slightly negative. First, we extends the result [LS16, Theorem 3.3] to the generalized equation (1.6) as follows:
Theorem 1.1.
Let $X$ be a compact complex manifold, and $T$, $v$, $\omega$, $\widehat{\chi}$ as above. Then the generalized equation (1.6) with $b\geqslant 0$ admits a solution if and only if there exists $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$ such that
$\big(c+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0.$ (1.8)
In the study of canonical Kähler metrics, a fundamental question is whether the solvability of an equation can be related to the properness or coercivity of a certain energy functional. For instance, Tian’s properness conjecture [Tia94] predicted that the existence of Kähler–Einstein metrics is equivalent to the properness of the $K$-energy. When the manifold has trivial automorphism group, Ding–Tian [DT92] proved one direction (properness $\Rightarrow$ existence), and Tian [Tia97a] confirms a converse direction (existence $\Rightarrow$ properness). In general case, Darvas–Rubinstein [DR17] proposed a refined version of Tian’s properness conjecture and proved it by exploiting a general framework in Finsler geometry. For constant scalar curvature Kähler metrics, a version of Tian’s properness conjecture was proved by Berman–Darvas–Lu [BDL20] and Chen–Cheng [CC18, CC21a, CC21b]. For complex Hessian equations, Collins–Székelyhidi [CS17] obtained a similar result for the $J$-equation, and Chu–Lee [CL21] for the hypercritical deformed Hermitian–Yang–Mills equation.
Motivated by above works, we show the equivalence between the existence and coercivity for the modified $J$-equation:
Theorem 1.2.
Let $X$ be a compact complex manifold, and $T$, $v$, $\omega$, $\widehat{\chi}$ as above. Then the modified $J$-equation (1.2) admits a solution $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ if and only if the modified $J$-functional $J_{v}^{\omega}$ is coercive.
For the precise definition of $J_{v}^{\omega}$ and coercivity, see Section 2.4. To prove the above theorem, we use the local smoothing and gluing method established in [Che21, CS17]. A new issue arising in the modified case is that even if we approximate the holomorphic vector field $v$ locally by vector fields with constant coefficients, to deal with the approximation error terms, we need a uniform gradient estimate for the potentials along the modified $J$-flow. To avoid this, we consider the product of $X$ with an annulus, add a rotational vector field $w$ to $v$ so as to eliminate the zeroes. We take holomorphic coordinates on which $v+w$ has constant coefficients by using the rectification theorem. Then we perform smoothing, gluing and finally averaging in the rotational direction to cancel out the contribution of $w$ (see Section 4.1.1 for more details).
In practice, it seems to be difficult to produce $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$ satisfying (1.5). In particular, it is not clear whether the existence of a solution to the equation depends only on the classes $\alpha$, $\beta$. The following theorem settles this question by showing that solvability depends only on the class $\alpha$.
Theorem 1.3.
Let $X$ be a compact complex manifold, and $T$, $v$, $\omega$, $\widehat{\chi}$ as above. Suppose that there exists a $T$-invariant Kähler form $\chi\in\beta$ satisfying (1.2) with respect to $\omega$ and $\omega^{\prime}\in\alpha$ is another Kähler form. Then there exists $\chi^{\prime}\in\beta$ that solves (1.2) with respect to $\omega^{\prime}$.
On the other hand, there is the question of whether the existence of solutions to the equation can be characterized numerically, or algebraically. For instance, the Nakai–Moishezon criterion [DP04] allows one to determine the Kähler condition in terms of intersection numbers with arbitrary subvarieties. Theorem 1.3 strongly indicates that a similar criterion should hold for the modified $J$-equation as well. More precisely, we would like to propose the following:
Conjecture 1.4.
Let $X$ be a compact complex manifold, and $T$, $v$, $\omega$, $\widehat{\chi}$ as above. Then there exists $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ satisfying (1.2) if and only if $m_{X}>0$ and for all $T$-invariant $p$-dimensional irreducible subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ we have
$\int_{Y}\big(\big(c_{X}+\theta_{v}^{X}(\chi)\big)\chi^{p}-p\omega\wedge\chi^{p-1}\big)>0,$ (1.9)
where $\chi\in\beta$ is any $T$-invariant Kähler form.
Remark 1.5.
The condition (1.9) is independent of the choice of $\chi\in\beta$. Indeed, the term including the Hamiltonian $\theta_{v}^{X}(\chi)$ can be rewritten as
$\int_{Y}\theta_{v}^{X}(\chi)\chi^{p}=\frac{1}{p+1}\int_{Y}(\chi+\theta_{v}^{X}(\chi))^{p+1}.$
Set $d_{v}:=d-4\pi i_{\mathrm{Im}(v)}$ so that $\chi+\theta_{v}^{X}(\chi)$ is $d_{v}$-closed. For another $T$-invariant Kähler form $\chi^{\prime}=\chi+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi$ with a $T$-invariant potential $\phi$, a direct computation shows that
$\chi^{\prime}+\theta_{v}^{X}(\chi^{\prime})=\chi+\theta_{v}^{X}(\chi)+\frac{1}{4\pi}d_{v}d^{c}\phi$
where $d^{c}:=\sqrt{-1}(\bar{\partial}-\partial)$. So the desired statement follows from the equivariant Chern–Weil theory (cf. [BGV92]).
[…]
Theorem 1.6.
Let $X$ be a smooth projective toric variety, $T$ the real torus associated with $X$, and $v$, $\omega$, $\widehat{\chi}$ as above. Then there exists a solution $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ to the modified $J$-equation (1.2) if and only if $m_{X}>0$ and for all $p$-dimensional irreducible toric subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ we have
$\int_{Y}\big(\big(c_{X}+\theta_{v}^{X}(\chi)\big)\chi^{p}-p\omega\wedge\chi^{p-1}\big)>0,$
where $\chi\in\beta$ is any $T$-invariant Kähler form.
When dealing with the contribution of the holomorphic vector field $v$, one need to handle carefully the fact that the normalization of the Hamiltonian (1.1) differs for each subvariety $Y$. Also, when gluing subsolutions, we need to control the derivatives of potential functions along $v$ near each $D_{i}$, using a $T$-equivariant linearization of $\mathcal{O}_{X}(D_{i})$. For this reason, we make the additional assumption that $X$ is projective. As an application, we also obtain:
Theorem 1.7.
Let $X$ be an $n$-dimensional smooth projective toric variety with real torus $T$, and let $\widehat{\chi}$ be a $T$-invariant Kähler form on $X$ with Kähler class $\beta$. Let $\delta_{\beta}^{T^{\mathbb{C}}}$ be the $T^{\mathbb{C}}$-reduced delta invariant of $\beta$ defined by (6.2). Let $v_{\rm ext}$ be the extremal vector field associated with $\beta$ such that $\mathrm{Im}(v_{\rm ext})\in\mathfrak{t}$. Assume that there exists a constant $\varepsilon>0$ satisfying the following properties:
  1. (1)
    The class $(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler.
  2. (2)
    For any $T$-invariant Kähler forms $\chi\in\beta$,
    $\max_{X}\theta_{v_{\rm ext}}^{X}(\chi)<n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}.$
  3. (3)
    For any $p$-dimensional irreducible toric subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ and any $T$-invariant Kähler forms $\chi\in\beta$,
    $\int_{Y}\bigg[\bigg((n-p)(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}-\theta_{v_{\rm ext}}^{X}(\chi)\bigg)\chi^{p}+p\chi^{p-1}\cdot c_{1}(X)\bigg]>0.$
Then the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric.
Remark 1.8.
We emphasize that the above theorem is essentially established in [DJ25, Theorem 1.4]. Our contribution is limited in the toric case. Specifically, in the course of the proof, we need to show the existence of a solution to a certain modified $J$-equation, which we reformulate in numerical terms by using Theorem 1.6. Thus, if Conjecture 1.4 is true, Theorem 1.7 still holds for arbitrary compact Kähler manifolds. Also, it should be noted that in the toric case, a stronger statement holds: the conclusions remain valid even if $\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon$ in assumptions (1)–(3) is replaced by $\delta_{\beta}^{T^{\mathbb{C}}}$. This is due to the fact that, in (3), the object under considerations are restricted to toric subvarieties, which form a finite set.
We note that in the case $v=0$, Theorem 1.2, Theorem 1.3 and Theorem 1.6 were previously established in [CS17]. Also, a variant of Theorem 1.7 was obtained in [Che21, Corollary 1.5].
The organization of this paper is as follows. In Section 2 we collect some fundamental properties that will be used throughout this paper. In Section 3 we establish the long-time existence of the modified $J$-flow. In particular we prove the convergence of the flow when $b\geqslant 0$ under the existence of solusolutions, which gives a proof of Theorem 1.1. In Section 4 we show that the existence of solutions is equivalent to the coercivity of the functional $J_{v}^{\omega}$ (Theorem 1.2). As a corollary, we also obtain a proof of Theorem 1.3. In Section 5 we establish a Nakai–Moishezon type criterion for smooth projective toric varieties, which provides the proof of Theorem 1.6. Finally, in Section 6, we give the proof of Theorem 1.7.
Acknowledgment.
This author was supported by JSPS KAKENHI Grant Numbers 20K14308, 24K06730. The author is also greatful to Vamsi Pritham Pingali for pointing out an error of the proof of Theorem 1.2 in the previous version.
2. Preliminaries
2.1. Convexity properties
Many of the results in this subsection can be found in [CS17, Section 2]. Let $\Gamma\subset\mathbb{R}^{n}$ be the positive orthant and
$S_{k}(\lambda):=\sum_{1\leqslant j_{1}<\ldots<j_{k}\leqslant n}\lambda_{j_{1}}\ldots\lambda_{j_{k}},\quad\lambda\in\Gamma$
the elementary symmetric function of degree $k$. We have $S_{0}(\lambda)=1$ and $S_{-1}(\lambda)=0$. For any constant $b\in\mathbb{R}$ we define
$f_{b}(\lambda):=\frac{S_{n-1}(\lambda)}{S_{n}(\lambda)}+b\frac{S_{0}(\lambda)}{S_{n}(\lambda)}=\sum_{i=1}^{n}\frac{1}{\lambda_{i}}+\frac{b}{\lambda_{1}\cdots\lambda_{n}},\quad\lambda\in\Gamma.$
Let $\mathcal{M}$ be the space of all $n\times n$ positive definite Hermitian matrices. Set
$F_{b}(A):=f_{b}(\lambda),$
where $\lambda=(\lambda_{1},\ldots,\lambda_{n})$ denotes the eigenvalues of $A$ (up to ordering). Then at a diagonal matrix $A\in\mathcal{M}$ with eigenvalues $\lambda_{i}$, the derivative of $F_{b}$ is given by
$\partial_{ij}F_{b}=\partial_{i}f_{b}\cdot\delta_{ij},\quad\partial_{ij}\partial_{rs}F_{b}=\partial_{i}\partial_{r}f_{b}\cdot\delta_{ij}\delta_{rs}+\frac{\partial_{i}f_{b}-\partial_{j}f_{b}}{\lambda_{i}-\lambda_{j}}(1-\delta_{ij})\delta_{ir}\delta_{js},$ (2.1)
where we will regard the quotient appeared in the last term as the limit when $\lambda_{i}=\lambda_{j}$ [And94, Ger96]. Also for $A\in\mathcal{M}$ we define
$P(A):=\max_{k=1,\ldots,n}\sum_{i\neq k}\frac{1}{\lambda_{k}},\quad Q(A):=F_{0}(A)=\sum_{i=1}^{n}\frac{1}{\lambda_{i}},\quad R(A):=\max_{I}\sum_{i\in I}\frac{1}{\lambda_{i}},$
where the maximum in $R(A)$ is taken over all subsets $I\subset\{1,\ldots,n\}$ satisfying |I|=n2|I|=n-2. For $K>0$ we define convex subsets $\Gamma_{K}:=\{\lambda\in\Gamma|f_{0}(\lambda)<K\}$ and $\mathcal{M}_{K}:=\{A\in\mathcal{M}|Q(A)<K\}$. Although $P$ and $R$ are not necessarily smooth except at points where there is a spectral gap, they can be approximated by smooth convex functions, and hence are convex.
Lemma 2.1.
The functions $P$ and $R$ are convex on $\mathcal{M}$.
Proof.
The proof is essentially the same as that of [CLT24, Proposition 2.2 (4)]. We define a continuous, symmetric, and convex function $g\colon\Gamma\to\mathbb{R}$ by $g(\lambda):=\max_{k=1,\ldots,n}\sum_{i\neq k}\frac{1}{\lambda_{i}}$. For any fixed $A,A^{\prime}\in\mathcal{M}$, we take $K>0$ so that $A,A^{\prime}\in\mathcal{M}_{K}$. Let us consider the convex subset $\Gamma_{K}\subset\Gamma$ corresponding to $\mathcal{M}_{K}$. If $\lambda\in\Gamma_{K}$, then $\lambda_{i}>K^{-1}$ holds for all $i$. Thus, for any $r\in(0,\frac{1}{2}K^{-1})$, we can define the smoothing $g^{(r)}$ of $g$ at scale $r$ by
$g^{(r)}(\lambda):=\int_{\mathbb{R}^{n}}r^{-n}\rho\bigg(\frac{|\mu|}{r}\bigg)g(\lambda-\mu)d\mu,\quad\lambda\in\Gamma_{K},$
where $\rho(t)$ is a smooth, non-negative function supported on $[0,1]$, constant on $[0,1/2]$, and normalized so that $\int_{B_{1}(0)}\rho(|\mu|)d\mu=1$. Then $g^{(r)}$ is a smooth function on $\Gamma_{K}$ and converges uniformly to $g$ on $\Gamma_{K}$ as $r\to 0$. For any $\lambda,\lambda^{\prime}\in\Gamma_{K}$ and $t\in[0,1]$, the convexity of $g$ yields that
$g^{(r)}(t\lambda+(1-t)\lambda^{\prime})$ $=\int_{\mathbb{R}^{n}}r^{-n}\rho\bigg(\frac{|\mu|}{r}\bigg)g(t\lambda+(1-t)\lambda^{\prime}-\mu)d\mu$
$=\int_{\mathbb{R}^{n}}r^{-n}\rho\bigg(\frac{|\mu|}{r}\bigg)g(t(\lambda-\mu)+(1-t)(\lambda^{\prime}-\mu))d\mu$
$\leqslant\int_{\mathbb{R}^{n}}r^{-n}\rho\bigg(\frac{|\mu|}{r}\bigg)\big(tg(\lambda-\mu)+(1-t)g(\lambda^{\prime}-\mu)\big)d\mu$
$=tg^{(r)}(\lambda)+(1-t)g^{(r)}(\lambda^{\prime}),$
which shows that $g^{(r)}$ is convex on $\Gamma_{K}$. Let $\mathfrak{S}_{n}$ be the symmetric group of degree $n$ which acts linearly on $\Gamma_{K}$ by $\sigma\cdot\lambda:=(\lambda_{\sigma(1)},\ldots,\lambda_{\sigma(n)})$. Then the symmetrization $g_{r}(\lambda):=\frac{1}{n!}\sum_{\sigma\in\mathfrak{S}_{n}}g^{(r)}(\sigma\cdot\lambda)$ is convex and converges uniformly to $g$ on $\Gamma_{K}$ as $r\to 0$. Therefore, we can apply the result [Spr05] to $g_{r}$, and deduce that the function $P_{r}$ on $\mathcal{M}_{K}$ corresponding to $g_{r}$ is convex. Thus we have
$P_{r}(tA+(1-t)A^{\prime})\leqslant tP_{r}(A)+(1-t)P_{r}(A^{\prime}),\quad t\in[0,1].$
Taking the limit $r\to 0$, we obtain
$P(tA+(1-t)A^{\prime})\leqslant tP(A)+(1-t)P(A^{\prime}),\quad t\in[0,1].$
Since $A$ and $A^{\prime}$ are arbitrarily chosen in $\mathcal{M}$, this shows that $P$ is convex on $\mathcal{M}$. The same argument applies to $R$ as well. ∎
For Kähler forms $\omega$, $\chi$ on a complex manifold $X$, we often write $g$, $h$ as the Riemannian metric corresponding to $\omega$, $\chi$ respectively, i.e.
$\omega=\frac{\sqrt{-1}}{2\pi}\sum_{i,j}g_{i\bar{j}}dz^{i}\wedge dz^{\bar{j}},\quad\chi=\frac{\sqrt{-1}}{2\pi}\sum_{i,j}h_{i\bar{j}}dz^{i}\wedge dz^{\bar{j}}$
in local coordinates. We define the matrix $A$ by $A_{j}^{i}:=g^{i\bar{k}}h_{j\bar{k}}$, and set
$F_{\omega,b}(\chi):=F_{b}(A),\quad P_{\omega}(\chi):=P(A),\quad Q_{\omega}(\chi):=Q(A),\quad R_{\omega}(\chi):=R(A),$
where the right hand side of each of the above expression is computed of the pointwise eigenvalues $\lambda_{i}$ of $A$. It is also worth nothing that the eigenvalues of $A$ do not depend on the choice of local coordinates $(z^{1},\ldots,z^{n})$, but only on $\omega$ and $\chi$. For a submanifold $Y\subset X$, we often consider the restriction of the operators $P_{\omega}$, $Q_{\omega}$ to $Y$:
$P_{Y,\omega}(\chi):=P_{\omega|_{Y}}(\chi|_{Y}),\quad Q_{Y,\omega}(\chi):=Q_{\omega|_{Y}}(\chi|_{Y}).$
Then we have the following:
Proposition 2.2.
  1. (1)
    For any submanifold $Y\subset X$ we have
    $Q_{Y,\omega}(\chi)\leqslant Q_{X,\omega}(\chi).$
  2. (2)
    For any hypersurface $Y\subset X$ we have
    $P_{Y,\omega}(\chi)\leqslant R_{X,\omega}(\chi).$
Proof.
This follows immediately from the Courant–Fischer–Weyl type min-max principle established in [Che21, Section 3] (see also [CLT24, Section 2]), but for the reader’s convenience, we provide a detailed proof here.
Let $p:=\dim Y$. At each point $x\in Y$, we choose an $\omega$-orthonormal basis $e_{1},\ldots,e_{n}$ of $T_{x}X$ and identify $T_{x}X\simeq\mathbb{C}^{n}$. Moreover, we choose the basis such that $e_{1},\ldots,e_{p}$ span $T_{x}Y$, so that we obtain $T_{x}Y\simeq\mathbb{C}^{p}\simeq\mathbb{C}^{p}\times\{0\}\subset\mathbb{C}^{n}$. This allows us to identify $\chi$ with an $n\times n$ positive definite Hermitian matrix $A$. We shall regard $A$ as a bilinear form on $\mathbb{C}^{n}$.
For any positive integer $k\leqslant p$ and $k$-dimensional subspace $V$ of $\mathbb{C}^{n}$ there exists $n\times k$ matrix $U\in\operatorname{Mat}(n,k;\mathbb{C})$ such that $\overline{U}^{T}U=I_{k}$ and a basis of $V$ consists of the columns of $U$. The representation of the restriction $A|_{V}$ with respect to the basis $U$ is then given by $\overline{U}^{T}AU$. In particular, we note that $\operatorname{Tr}(A|_{V})^{-1}:=\operatorname{Tr}(\overline{U}^{T}AU)^{-1}$ is independent of the choice of $U$ and depends only on $A$ and $V$. Also if $V\subset\mathbb{C}^{p}$ then $U$ can be decomposed as
$U=\begin{pmatrix}U^{\prime}\\ 0\end{pmatrix}$
for a matrix $U^{\prime}\in\operatorname{Mat}(p,k;\mathbb{C})$ such that $\overline{U^{\prime}}^{T}U^{\prime}=I_{k}$. Thus denoting by $A^{\prime}$ the $p\times p$ upper-left block of $A$, we obtain $\overline{U}^{T}AU=\overline{U^{\prime}}^{T}A^{\prime}U^{\prime}$, which shows that the definition of $\operatorname{Tr}(A|_{V})^{-1}$ is compatible with restriction to $\mathbb{C}^{p}$.
(1) We choose $U$ such that $\overline{U}^{T}AU=\operatorname{diag}(\lambda_{1}^{\prime},\ldots,\lambda_{p}^{\prime})$ and $S\in U(n)$ such that $S_{i\bar{j}}=U_{i\bar{j}}$ ($i=1,\ldots,n$; $j=1,\ldots,p$). Then by using the Schur–Horn theorem we know that the diagonal $(\lambda_{1}^{\prime},\ldots,\lambda_{p}^{\prime},(\overline{S}^{T}AS)_{p+1,\overline{p+1}},\ldots,(\overline{S}^{T}AS)_{n,\overline{n}})$ of the matrix $\overline{S}^{T}AS$ lies in the convex full of the vectors obtained by permuting the entries of $(\lambda_{1},\ldots,\lambda_{n})$, where $\lambda_{i}$ are eigenvalues of $A$. Let $I\subset\{1,\ldots,n\}$ be a subset with $|I|=p$. Combining with the convexity of the function $\Gamma\ni(\mu_{1},\ldots,\mu_{n})\mapsto\max_{I}\sum_{i\in I}\frac{1}{\mu_{i}}\in\mathbb{R}$ we know that
$Q_{X,\omega}(\chi)$ $=\sum_{i=1}^{n}\frac{1}{\lambda_{i}}\geqslant\max_{I}\sum_{i\in I}\frac{1}{\lambda_{i}}=\max_{\begin{subarray}{c}U\in\operatorname{Mat}(n,p;\mathbb{C});\\ \overline{U}^{T}U=I_{p},\;\overline{U}^{T}AU=\operatorname{diag}(\lambda_{1}^{\prime},\ldots,\lambda_{p}^{\prime})\end{subarray}}\sum_{i=1}^{p}\frac{1}{\lambda_{i}^{\prime}}=\max_{V\subset\mathbb{C}^{n}}\operatorname{Tr}(A|_{V})^{-1}$
$\geqslant\operatorname{Tr}(A|_{\mathbb{C}^{p}})^{-1}=Q_{Y,\omega}(\chi),$
where the maximum is taken over all $p$-dimensional subspaces $V\subset\mathbb{C}^{n}$.
(2) As in (1), we can deduce that
$R_{X,\omega}(\chi)=\max_{V\subset\mathbb{C}^{n}}\operatorname{Tr}(A|_{V})^{-1}\geqslant\max_{V\subset\mathbb{C}^{n-1}}\operatorname{Tr}(A|_{V})^{-1}=P_{Y,\omega}(\chi),$
where the maxima are taken over all (n2)(n-2)-dimensional subspaces $V\subset\mathbb{C}^{n}$ or $V\subset\mathbb{C}^{n-1}$. ∎
2.1.1. The $b=-\varepsilon<0$ case
Next we consider the case $b=-\varepsilon<0$.
Proposition 2.3.
For any $K>0$ there exists $\varepsilon_{1}=\varepsilon_{1}(n,K)>0$ such that for all $\varepsilon\in(0,\varepsilon_{1})$, the function $f_{-\varepsilon}\colon\Gamma_{K}\to\mathbb{R}$ satisfies the following properties:
  1. (1)
    $f_{-\varepsilon}>0$.
  2. (2)
    $\partial_{i}f_{-\varepsilon}<0$ for all $i$.
  3. (3)
    If $\lambda_{i}\geqslant\lambda_{j}$, then $\partial_{i}f_{-\varepsilon}\geqslant\partial_{j}f_{-\varepsilon}$.
  4. (4)
    $f_{-\varepsilon}$ is convex.
Proof.
For notational convenience let
$S_{n-1;i}(\lambda):=\frac{S_{n}(\lambda)}{\lambda_{i}}.$
Since $\lambda_{i}>K^{-1}$ from the condition $f_{0}(\lambda)<K$ we have
$f_{-\varepsilon}(\lambda)=\frac{S_{n-1}(\lambda)-\varepsilon}{S_{n}(\lambda)}\geqslant\frac{(n-1)K^{-n+1}-\varepsilon}{S_{n}(\lambda)},$
$\partial_{i}f_{-\varepsilon}(\lambda)=\frac{-S_{n-1;i}(\lambda)+\varepsilon}{\lambda_{i}S_{n}(\lambda)}\leqslant\frac{-K^{-n+1}+\varepsilon}{\lambda_{i}S_{n}(\lambda)}.$
Also if $\lambda_{i}\geqslant\lambda_{j}$ then
$\partial_{i}f_{-\varepsilon}(\lambda)-\partial_{j}f_{-\varepsilon}(\lambda)$ $=\frac{(\lambda_{i}-\lambda_{j})(S_{n-1;i}(\lambda)+S_{n-1;j}(\lambda)-\varepsilon)}{\lambda_{i}\lambda_{j}S_{n}(\lambda)}$
$\geqslant\frac{(\lambda_{i}-\lambda_{j})(2K^{-n+1}-\varepsilon)}{\lambda_{i}\lambda_{j}S_{n}(\lambda)}.$
Finally, we compute
$\partial_{i}\partial_{j}f_{-\varepsilon}(\lambda)=\frac{2S_{n-1;i}(\lambda)\delta_{ij}-\varepsilon(1+\delta_{ij})}{\lambda_{i}\lambda_{j}S_{n}(\lambda)}.$
Thus for all $w=(w_{1},\ldots,w_{n})\in\mathbb{R}^{n}$ we have
$\sum_{i,j}w_{i}w_{j}\partial_{i}\partial_{j}f_{-\varepsilon}(\lambda)$ $=\frac{2}{S_{n}(\lambda)}\sum_{i}\frac{w_{i}^{2}}{\lambda_{i}^{2}}S_{n-1;i}(\lambda)-\frac{\varepsilon}{S_{n}(\lambda)}\bigg(\sum_{i}\frac{w_{i}}{\lambda_{i}}\bigg)^{2}-\frac{\varepsilon}{S_{n}(\lambda)}\sum_{i}\frac{w_{i}^{2}}{\lambda_{i}^{2}}$
$\geqslant\frac{2K^{-n+1}-\varepsilon(n+1)}{S_{n}(\lambda)}\sum_{i}\frac{w_{i}^{2}}{\lambda_{i}^{2}}.$
The above computations show that all of the required statements hold if $\varepsilon$ is sufficiently small. ∎
Thus combining with (2.1), we know that $F_{-\varepsilon}\colon\mathcal{M}_{K}\to\mathbb{R}$ is convex for $\varepsilon\in(0,\varepsilon_{1})$ (cf. [And94, Ger96, Spr05]). However, we need a stronger convexity property as follows:
Proposition 2.4 ([CS17], Lemma 9).
For any $K>0$ there exists $\varepsilon_{2}(n,K)>0$ such that for all $\varepsilon\in(0,\varepsilon_{2})$, diagonal matrix $A\in\mathcal{M}_{K}$ with entries $\lambda_{i}$ and Hermitian matrix $B_{ij}$ we have
$\sum_{p,q,r,s}B_{rs}\overline{B_{qp}}\partial_{pq}\partial_{rs}F_{-\varepsilon}(A)+\sum_{i,j}|B_{ij}|^{2}\frac{\partial_{ii}F_{-\varepsilon}(A)}{\lambda_{j}}\geqslant 0.$
We also need the following estimate to deal with the modified $J$-flow with negative twist $b=-\varepsilon<0$.
Lemma 2.5.
For any $K>0$ and $C>0$ there exists a constant $\varepsilon_{3}=\varepsilon_{3}(n,K,C)>0$ such that for any $\varepsilon\in(0,\varepsilon_{3})$, we have the following: if the continuous path $A_{t}\in\mathcal{M}$ ($t\in[0,1]$) satisfies
$Q(A_{0})<K,\quad F_{-\varepsilon}(A_{t})<K+2C\quad(t\in[0,1]),$
then $Q(A_{t})<K+3C$ for $t\in[0,1]$.
Proof.
We follow closely to the argument [CS17, Lemma 19]. Let $\lambda_{i}$ be the eigenvalue of $A_{t}$. Assume that $Q(A_{t})=K+3C$ for some $t\in(0,1]$. Then AM-GM inequality shows that
$\frac{1}{\lambda_{1}\cdots\lambda_{n}}\leqslant n^{-n}\bigg(\sum_{i=1}^{n}\frac{1}{\lambda_{i}}\bigg)^{n}=n^{-n}(K+3C)^{n}.$
If we set $\varepsilon_{3}:=\frac{n^{n}C}{2(K+3C)^{n}}$ then for any $\varepsilon\in(0,\varepsilon_{3})$ we have
$F_{-\varepsilon}(A_{t})=\sum_{i=1}^{n}\frac{1}{\lambda_{i}}-\frac{\varepsilon}{\lambda_{1}\cdots\lambda_{n}}\geqslant K+3C-\varepsilon n^{-n}(K+2C)^{n}>K+2C$
which contradicts the assumption. So we get $Q(A_{t})<K+3C$ for all $t\in[0,1]$. ∎
Lemma 2.6.
Let $\omega_{1}$, $\omega_{2}$, $\chi$ be Kähler forms on $X$. Assume that $Q_{\omega_{2}}(\chi)\leqslant K$ and
$(1-\sigma)\omega_{1}\leqslant\omega_{2}\leqslant(1+\sigma)\omega_{1}$
for some $K>0$ and $0<\sigma<1/2$. Then we have
$|Q_{\omega_{1}}(\chi)-Q_{\omega_{2}}(\chi)|\leqslant 2K\sigma,\quad|R_{\omega_{1}}(\chi)-R_{\omega_{2}}(\chi)|\leqslant 2K\sigma.$
Proof.
Let $\lambda_{1}\leqslant\ldots\leqslant\lambda_{n}$ and $\mu_{1}\leqslant\ldots\leqslant\mu_{n}$ be eigenvalues of $\chi$ with respect to $\omega_{1}$ and $\omega_{2}$ respectively. Then the assumption yields that
$(1-\sigma)\mu_{i}\leqslant\lambda_{i}\leqslant(1+\sigma)\mu_{i},\quad i=1,\ldots,n.$
From the assumption $Q_{\omega_{2}}(\chi)\leqslant K$ we get
$|Q_{\omega_{1}}(\chi)-Q_{\omega_{2}}(\chi)|\leqslant\sum_{i=1}^{n}\bigg|\frac{1}{\lambda_{i}}-\frac{1}{\mu_{i}}\bigg|=\sum_{i=1}^{n}\frac{|\lambda_{i}-\mu_{i}|}{\lambda_{i}\mu_{i}}\leqslant\frac{\sigma}{1-\sigma}\sum_{i=1}^{n}\frac{1}{\mu_{i}}\leqslant 2K\sigma.$
The proof for $R$ is similar. ∎
Lemma 2.7.
Let $\omega$ and $\chi$ be Kähler forms and $K>0$ a constant. Assume that $Q_{\omega}(\chi)\leqslant K$. Then there exists $\varepsilon_{4}=\varepsilon_{4}(n,K)>0$ such that for all $\varepsilon\in(0,\varepsilon_{4})$ we have
$P_{\omega}(\chi)\leqslant F_{\omega,-\varepsilon}(\chi).$
Proof.
Let $\lambda_{1}\leqslant\ldots\leqslant\lambda_{n}$ be the eigenvalues of $\chi$ with respect to $\omega$. By using $\lambda_{i}\geqslant K^{-1}$ we compute
$F_{\omega,-\varepsilon}(\chi)=P_{\omega}(\chi)+\frac{1}{\lambda_{n}}\bigg(1-\frac{\varepsilon}{\lambda_{1}\cdots\lambda_{n-1}}\bigg)\geqslant P_{\omega}(\chi)+\frac{1}{\lambda_{n}}(1-\varepsilon K^{n-1}).$
So if we set $\varepsilon_{4}=K^{-n+1}$ then for any $\varepsilon\in(0,\varepsilon_{4})$, we have
$P_{\omega}(\chi)\leqslant F_{\omega,-\varepsilon}(\chi).$
2.1.2. The $b\geqslant 0$ case
Now we consider the case $b\geqslant 0$. In this case, the situation is considerably easier to handle compared to the case $b=-\varepsilon<0$.
Proposition 2.8.
For $b\geqslant 0$ the function $f_{b}\colon\Gamma\to\mathbb{R}$ satisfies the following properties:
  1. (1)
    $f_{b}>0$.
  2. (2)
    $\partial_{i}f_{b}<0$ for all $i$.
  3. (3)
    If $\lambda_{i}\geqslant\lambda_{j}$, then $\partial_{i}f_{b}\geqslant\partial_{j}f_{b}$.
  4. (4)
    $f_{b}$ is convex.
Proof.
The calculation in Proposition 2.3 can be applied verbatim to this case. ∎
Thus the propeties (3), (4) together with (2.1) implies that $F_{b}\colon\mathcal{M}\to\mathbb{R}$ is convex. A strong convexity property holds in the same manner:
Proposition 2.9 ([CS17], Lemma 8).
For any diagonal matrix $A\in\mathcal{M}$ with eigenvalues $\lambda_{i}$ and Hermitian matrix $B_{ij}$ we have
$\sum_{p,q,r,s}B_{rs}\overline{B_{qp}}\partial_{pq}\partial_{rs}F_{b}(A)+\sum_{i,j}|B_{ij}|^{2}\frac{\partial_{ii}F_{b}(A)}{\lambda_{j}}\geqslant 0.$
2.2. Hamiltonians
Let $X$ be a compact complex manifold, $T\subset\operatorname{Aut}_{\rm red}(X)$ a real torus, $\widehat{\chi}$ a $T$-invariant Kähler form on $X$, and $v$ a holomorphic vector field satisfying $\mathrm{Im}(v)\in\mathfrak{t}$.
Lemma 2.10.
For any $\chi_{\phi}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi$ with $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, the following statement holds:
  1. (1)
    The function $\theta_{v}^{X}(\chi_{\phi})$ is $T$-invariant.
  2. (2)
    We have $\theta_{v}^{X}(\chi_{\phi})=\theta_{v}^{X}(\widehat{\chi})+v(\phi)$. In particular, $\max_{X}\theta_{v}(\chi_{\phi})$ as well as $\min_{X}\theta_{v}(\chi_{\phi})$ is independent of the choice of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$.
Proof.
(1) By (1.1), we have
$i_{\mathrm{Im}(v)}\chi_{\phi}=\frac{1}{4\pi}d\theta_{v}^{X}(\chi_{\phi}),$
which shows that $d\theta_{v}^{X}(\chi_{\phi})$ is $T$-invariant. So for any $\tau\in T$ we have $d\tau^{\ast}\theta_{v}^{X}(\chi_{\phi})=\tau^{\ast}d\theta_{v}^{X}(\chi_{\phi})=d\theta_{v}^{X}(\chi_{\phi})$, from which it follows that $\tau^{\ast}\theta_{v}^{X}(\chi_{\phi})-\theta_{v}^{X}(\chi_{\phi})=C$ for some constant $C\in\mathbb{R}$. By using (1.1) and the fact that $\chi_{\phi}$ is $T$-invariant, integrating over $X$ with respect to $\chi_{\phi}^{n}$ yields $C=0$. Hence $\theta_{v}^{X}(\chi_{\phi})$ is $T$-invariant.
(2) The result is obtained in [Zhu00, Corollary 5.3]. Applying $i_{v}$ to $\chi_{\phi}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi$ we observe that $\theta_{v}^{X}(\chi_{\phi})=\theta_{v}^{X}(\widehat{\chi})+v(\phi)$ modulo additive constants. Thus
$\theta_{v}^{X}(\chi_{t\phi})=\theta_{v}^{X}(\widehat{\chi})+tv(\phi)+C_{t},\quad t\in[0,1]$
for some $t$-dependent constant $C_{t}$. Differentiating $\int_{X}\theta_{v}^{X}(\chi_{t\phi})\chi_{t\phi}^{n}=0$ in $t$ and integrating by parts yield that
$0=\frac{d}{dt}\int_{X}\theta_{v}^{X}(\chi_{t\phi})\chi_{t\phi}^{n}=\int_{X}\big(v(\phi)+\frac{d}{dt}C_{t}+\theta_{v}^{X}(\chi_{t\phi})\Delta_{t\phi}\phi\big)\chi_{t\phi}^{n}=\beta^{n}\frac{d}{dt}C_{t}.$
Thus we have $C_{t}=C_{0}=0$ which shows the first assertion. For the second assertion we take a point $x_{0}\in X$ where $\theta_{v}(\chi_{\phi})$ achieves the maximum. Then $\bar{\partial}\theta_{v}(\chi_{\phi})(x_{0})=0$, so the vector $v$ vanishes at $x_{0}$. Combining with $\theta_{v}^{X}(\chi_{\phi})=\theta_{v}^{X}(\widehat{\chi})+v(\phi)$ we observe that
[…]
From the above lemma we know that the constant
$m_{X}=c_{X}+\min_{X}\theta_{v}^{X}(\chi_{\phi})$
is independent of a choice of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$. We fix a uniform constant $C_{\theta}>0$ (independent of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$) such that
$|\theta_{v}^{X}(\chi_{\phi})|\leqslant C_{\theta}.$ (2.2)
2.3. Subsolutions
In this subsection we study general properties for subsolutions (1.5) related to the regularized maximum and averaging over the torus $T$. We first discuss how the subsolution coindition can be expressed in terms of the operator $P_{\omega}$.
Proposition 2.11.
The subsolution condition (1.8) for $\chi:=\chi_{\underline{\phi}}$ is equivalent to
$P_{\omega}(\chi)-\theta_{v}^{X}(\chi)<c.$
In this case, we have
$\big(c+\theta_{v}^{X}(\chi)\big)\chi^{p}-p\omega\wedge\chi^{p-1}>0$ (2.3)
for all $p=1,\ldots,n-1$.
Proof.
This follows immediately from the computation in [Che21, Section 1]. We first note that the positivity condition (2.3) (in the sense of [Dem12, Chapter III, Section 1.A]) means that
$\big(c+\theta_{v}^{X}(\chi)\big)\chi^{p}-p\omega\wedge\chi^{p-1}\wedge\Omega>0$
for any simple positive (np,np)(n-p,n-p)-form $\Omega$. We choose a coordinates such that $g_{i\bar{j}}=\delta_{ij}$ and $h_{i\bar{j}}=\lambda_{i}\delta_{ij}$. Then
$\big(c+\theta_{v}^{X}(\chi)\big)\chi^{p}-p\omega\wedge\chi^{p-1}=p!\bigg(\frac{\sqrt{-1}}{2\pi}\bigg)^{p}\sum_{I}\bigg(c+\theta_{v}^{X}(\chi)-\sum_{i\in I}\frac{1}{\lambda_{i}}\bigg)\bigg(\prod_{i\in I}\lambda_{i}\bigg)dz^{I}\wedge d\bar{z}^{I},$
where the sum is taken over all subsets $I\subset\{1,\ldots,n\}$ with $|I|=p$. The form $\Omega$ can be expressed as
$\Omega=\bigg(\frac{\sqrt{-1}}{2\pi}\bigg)^{n-p}\sum_{I}a_{I}dz^{I^{c}}\wedge d\bar{z}^{I^{c}}+\Omega^{\prime},$
where $I^{c}$ denotes the complement of $I\subset\{1,\ldots,n\}$ and $\Omega^{\prime}$ is a smooth form such that $\Omega^{\prime}\wedge dz^{I}\wedge d\bar{z}^{I}=0$ for all $I$. Since $\Omega$ is a simple positive form, we have $a_{I}\geqslant 0$ for all $I$ and at least one $a_{I}$ is positive. Then a direct computation shows that
$\frac{\big(c+\theta_{v}^{X}(\chi)\big)\chi^{p}-p\omega\wedge\chi^{p-1}\wedge\Omega}{\omega^{n}}=\frac{p!}{n!}\sum_{I}a_{I}\bigg(c+\theta_{v}^{X}(\chi)-\sum_{i\in I}\frac{1}{\lambda_{i}}\bigg)\bigg(\prod_{i\in I}\lambda_{i}\bigg).$
The desired statement follows from the above formula. ∎
Following [Dem12, Section 5.E] for arbitrary $\eta=(\eta_{1},\ldots,\eta_{N})\in(0,\infty)^{N}$, we define the function $M_{\eta}\colon\mathbb{R}^{N}\to\mathbb{R}$ by
$M_{\eta}(t_{1},\ldots,t_{N}):=\int_{\mathbb{R}^{N}}\max\{t_{1}+h_{1},\ldots,t_{N}+h_{N}\}\prod_{j=1,\ldots,N}\theta\bigg(\frac{h_{j}}{\eta_{j}}\bigg)dh_{1}\ldots dh_{N},$
where $\theta$ is a non-negative smooth function on $\mathbb{R}$, supported on [1,1][-1,1], such that $\int_{\mathbb{R}}\theta(h)dh=1$ and $\int_{\mathbb{R}}h\theta(h)dh=0$. The fundamental properties for the regularized maximum are summed up as follows:
Lemma 2.12 ([Dem12], Lemma 5.18).
The function $M_{\eta}$ satisfies the following:
  1. (1)
    $M_{\eta}(t_{1},\ldots,t_{N})$ is non decreasing in all variables and convex on $\mathbb{R}^{N}$.
  2. (2)
    $\max\{t_{1},\ldots,t_{N}\}\leqslant M_{\eta}(t_{1},\ldots,t_{N})\leqslant\max\{t_{1}+\eta_{1},\ldots,t_{N}+\eta_{N}\}$.
  3. (3)
    $M_{\eta}(t_{1},\ldots,t_{N})=M_{(\eta_{1},\ldots,\check{\eta_{j}},\ldots,\eta_{N})}(t_{1},\ldots,\check{t_{j}},\ldots,t_{N})$ if $t_{j}+\eta_{j}\leqslant\max_{k\neq j}\{t_{k}-\eta_{k}\}$, where $\check{f}$ means eliminating $f$.
  4. (4)
    $M_{\eta}(t_{1}+a,\ldots,t_{N}+a)=M_{\eta}(t_{1},\ldots,t_{N})+a$ for all $a\in\mathbb{R}$.
By the property (1) and differentiating (4) at $a=0$ we have
$\sum_{j=1}\frac{\partial M_{\eta}}{\partial t_{j}}=1,\quad\frac{\partial M_{\eta}}{\partial t_{j}}\geqslant 0.$ (2.4)
Now let $X$ be a complex manifold, and let $\omega$ and $\widehat{\chi}$ be Kähler forms on $X$. Let $\{B_{j}\}_{j\in\{1,\ldots,N\}}$ be a finite covering of $X$ and $\varphi_{j}$ smooth functions on $B_{j}$ satisfying $\varphi_{j}(x)<\max_{k=1,\ldots,N}\{\varphi_{k}(x)\}$ at every point $x\in\partial B_{j}$ (where we extend $\varphi_{j}$ on $X$ so that $\varphi_{j}=$-$\infty$ outside $B_{j}$). In particular, we can choose a sufficiently small vector $\eta=(\eta_{1},\ldots,\eta_{N})\in\mathbb{R}^{N}$ so that $\varphi_{j}(x)+\eta_{j}\leqslant\max_{k=1,\ldots,N}\{\varphi_{k}(x)-\eta_{k}\}$ holds for all $j$ and $x\in\partial B_{j}$. Then as a corollary of the above lemma (see also [Dem12, Corollary 5.19]) we find that the function $\varphi:=M_{\eta}(\varphi_{1},\ldots,\varphi_{N})$ is smooth on $X$. Moreover, a direct computation shows that the second derivatives in local coordinates $(z^{1},\ldots,z^{n})$ are given by
$\frac{\partial^{2}\varphi}{\partial z^{k}\partial z^{\bar{\ell}}}=\sum_{a,b}\frac{\partial^{2}M_{\eta}}{\partial t_{a}\partial t_{b}}(\varphi_{1},\ldots,\varphi_{N})\cdot\frac{\partial\varphi_{a}}{\partial z^{k}}\cdot\frac{\partial\varphi_{b}}{\partial z^{\bar{\ell}}}+\sum_{a}\frac{\partial M_{\eta}}{\partial t_{a}}(\varphi_{1},\ldots,\varphi_{N})\cdot\frac{\partial^{2}\varphi_{a}}{\partial z^{k}\partial z^{\bar{\ell}}}.$ (2.5)
The following extends the argument in [Che21, Section 4]:
Proposition 2.13.
In the above setting, let $f\in C^{\infty}(X;\mathbb{R})$ and $w$ a holomorphic vector field on $X$. We further assume that $\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}>0$ on each $B_{j}$. Then for a sufficiently small vector $\eta\in\mathbb{R}^{N}$ chosen as above, the regularized maximum $\varphi=M_{\eta}(\varphi_{1},\ldots,\varphi_{N})$ satisfies $\chi:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi>0$. Moreover, if $\varphi_{j}$ satisfies
$P_{\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\bigg)-\mathrm{Re}(w)(\varphi_{j})<f$
on each $B_{j}$, then
$P_{\omega}(\chi)-\mathrm{Re}(w)(\varphi)<f.$
A similar formula holds when $P_{\omega}$ is replaced by $Q_{\omega}$ or $R_{\omega}$.
Proof.
By the convexity of $M_{\eta}$, (2.4) and (2.5) we observe that
$\chi\geqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j})>0.$
The term $\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot\big(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\big)$ is a weighted average of $\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}$ with $\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}=1$. Thus the monotonicity and convexity property of $P_{\omega}$ shows that
$P_{\omega}(\chi)\leqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot P_{\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\bigg),$
and hence
$P_{\omega}(\chi)-\mathrm{Re}(w)(\varphi)\leqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot\bigg[P_{\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\bigg)-\mathrm{Re}(w)(\varphi_{j})\bigg]<f.$
The above proof only requires the monotonicity and convexity of the operator $P_{\omega}$. So the same proof also works for $Q_{\omega}$ and $R_{\omega}$. ∎
In general, the Kähler form $\chi$ obtained in Proposition 2.13 is not necessarily $T$-invariant. For this reason we need the averaging argument as follows:
Proposition 2.14.
Let $X$ be a complex manifold, $T\subset\operatorname{Aut}(X)$ a real torus, and let $\omega$ and $\widehat{\chi}$ be $T$-invariant Kähler forms on $X$. Let $f\in C^{\infty}(X;\mathbb{R})^{T}$ and $w$ a holomorphic vector field on $X$ such that $\mathrm{Re}(w)$ is $T$-invariant. For a $\widehat{\chi}-$Kähler potential $\varphi$ on $X$, set $\chi:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi>0$,
$\widetilde{\varphi}:=\int_{\tau\in T}\tau^{\ast}\varphi dT,\quad\widetilde{\chi}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\widetilde{\varphi}$
so that $\widetilde{\chi}=\int_{\tau\in T}\tau^{\ast}\chi dT>0$, where $dT$ denotes the Haar measure on $T$ normalized by $\int_{T}dT=1$. If
$P_{\omega}(\chi)-\mathrm{Re}(w)(\varphi)<f,$
then the $T$-invariant potential $\widetilde{\varphi}$ satisfies
$P_{\omega}(\widetilde{\chi})-\mathrm{Re}(w)(\widetilde{\varphi})<f.$
A similar formula holds when $P_{\omega}$ is replaced by $Q_{\omega}$ or $R_{\omega}$.
Proof.
Since $\omega$ is $T$-invariant, the Kähler form $\tau^{\ast}\chi$ can be identified with a positive definite Hermitian matrix by using an $\omega$-orthonormal basis at each point on $X$. So the convexity of $P_{\omega}$ implies that
$P_{\omega}(\widetilde{\chi})\leqslant\int_{\tau\in T}\tau^{\ast}P_{\omega}(\chi)dT.$
Since $\mathrm{Re}(w)$ and $f$ are $T$-invariant, we obtain
$P_{\omega}(\widetilde{\chi})-\mathrm{Re}(w)(\widetilde{\varphi})\leqslant\int_{\tau\in T}\tau^{\ast}\big(P_{\omega}(\chi)-\mathrm{Re}(w)(\varphi)\big)dT<\int_{\tau\in T}\tau^{\ast}fdT=f.$
In the above proof, we only used the facts that $\omega$ is $T$-invariant and $P_{\omega}$ is convex. Hence the same assertion also holds for $Q_{\omega}$ and $R_{\omega}$. ∎
Proposition 2.15.
Let $X$ be a compact complex manifold, $T\subset\operatorname{Aut}_{\rm red}(X)$ a real torus, and let $\omega$ and $\widehat{\chi}$ be $T$-invariant Kähler forms on $X$. Let $v$ be a holomorphic vector field on $X$ satisfying $\mathrm{Im}(v)\in\mathfrak{t}$. For a $\widehat{\chi}-$Kähler potential $\varphi$ on $X$, set $\chi:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi>0$. If
$P_{\omega}(\chi)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\varphi)<c$
for some constant $c>0$, then the average $\widetilde{\chi}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\widetilde{\varphi}$ with $\widetilde{\varphi}:=\int_{\tau\in T}\tau^{\ast}\varphi dT$ satisfies
$P_{\omega}(\widetilde{\chi})-\theta_{v}^{X}(\widetilde{\chi})<c.$
A similar formula holds when $P_{\omega}$ is replaced by $Q_{\omega}$.
Proof.
[…]
2.4. Functionals on the space of Kähler potentials
Let $X$ be a compact complex manifold, $T\subset\operatorname{Aut}_{\rm red}(X)$ a real torus, and let $\omega\in\alpha$ and $\widehat{\chi}\in\beta$ be $T$-invariant Kähler forms. Define Aubin’s $I$-functional [Aub84] by the formula
$I(\phi):=\int_{X}\phi(\widehat{\chi}^{n}-\chi_{\phi}^{n}).$
Integrating by parts one can see that
$I(\phi)$ $=-\int_{X}\phi\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi\wedge(\chi_{\phi}^{n-1}+\chi_{\phi}^{n-2}\wedge\widehat{\chi}+\cdots+\widehat{\chi}^{n-1})$ (2.6)
$=\frac{\sqrt{-1}}{2\pi}\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge(\chi_{\phi}^{n-1}+\chi_{\phi}^{n-2}\wedge\widehat{\chi}+\cdots+\widehat{\chi}^{n-1})$
$\geqslant\int_{X}|\partial\phi|_{\widehat{h}}^{2}\widehat{\chi}^{n},$
where $\widehat{h}$ denotes the Riemannian metric corresponding to $\widehat{\chi}$. Let $b\in\mathbb{R}$, $c>0$ be constants related by (1.7). We define the modified $J$-functional by
$J_{v,b}^{\omega}(\phi):=\int_{0}^{1}\int_{X}\phi\big(F_{\omega,b}(\chi_{t\phi})-c-\theta_{v}^{X}(\chi_{t\phi})\big)\chi_{t\phi}^{n}dt,\quad\phi\in\mathcal{H}(X,\widehat{\chi})^{T}.$ (2.7)
By the definition the critical points of $J_{v,b}^{\omega}$ are just given by solutions to the generalized equation (1.6). When $b=0$, the functional $J_{v,b}^{\omega}$ was introduced by Li–Shi [LS16], which we will denote by $J_{v}^{\omega}$ for simplicity. Furthermore, when $v=0$, we shall abbreviate $J_{0}^{\omega}$ as $J^{\omega}$.
Proposition 2.16.
For $b\geqslant 0$, the functional $J_{v,b}^{\omega}$ is strictly convex along any $C^{1,1}$ geodesics in $\mathcal{H}(X,\widehat{\chi})^{T}$.
Proof.
Indeed, we can decompose the functional $J_{v,b}^{\omega}$ as
$J_{v,b}^{\omega}=J_{v}^{\omega}+b\int_{X}\phi\omega^{n}.$
The first term $J_{v}^{\omega}$ is strictly convex by [LS16, Proposition 3.1]. For $C^{2}$ geodesics $\frac{d^{2}}{dt^{2}}\phi_{t}-|\nabla_{t}\frac{d}{dt}\phi_{t}|_{h}^{2}=0$, we can easily see that
$\frac{d^{2}}{dt^{2}}\int_{X}\phi_{t}\omega^{n}=\int_{X}\big|\nabla_{t}\frac{d}{dt}\phi_{t}\big|_{h}^{2}\omega^{n}\geqslant 0.$
Then we may approximate $C^{1,1}$ geodesics by $\varepsilon$-geodesics just as in [Che00, Proposition 2.1]. ∎
Definition 2.17.
We say that the functional $J_{v,b}^{\omega}$ is coercive if there exist constants $\delta_{c}>0$, $B_{c}>0$ such that
$J_{v,b}^{\omega}(\phi)\geqslant\delta_{c}I(\phi)-B_{c}$
for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$.
Remark 2.18.
We define Aubin’s $J$-functional [Aub84] by
$J(\phi)=\int_{0}^{1}\int_{X}\phi(\widehat{\chi}^{n}-\chi_{t\phi}^{n})dt.$
Then it is well-known that
$\frac{1}{n+1}I(\phi)\leqslant J(\phi)\leqslant\frac{n}{n+1}I(\phi).$
Moreover, by a straightforward computation, one can see that
$J^{\widehat{\chi}}(\phi)=I(\phi)-J(\phi).$
In conclusion, we see that in the definition of coercivity, the functionals $I$, $J$, or $J^{\widehat{\chi}}$ may be used interchangeably. Alternatively, one can characterize coercivity equivariantly using the $d_{1}$-distance on $\mathcal{H}(X,\widehat{\chi})^{T}$ (for instance, see [DR17]).
3. The modified $J$-flow
Let $X$ be a compact complex manifold, $T\subset\operatorname{Aut}_{\rm red}(X)$ a real torus, and let $\omega\in\alpha$ and $\widehat{\chi}\in\beta$ be $T$-invariant Kähler forms on $X$. We define the modified $J$-flow as the gradient flow of (2.7), i.e.
$\frac{d}{dt}\phi_{t}=-F_{b}(A_{t})+c+\theta_{v}^{X}(\chi_{t}),$ (3.1)
where we set $A_{t}:=g^{i\bar{k}}h_{j\bar{k}}$ and $\chi_{t}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{t}$ for simplify the notations. Note that $\phi_{t}\in\mathcal{H}(X,\widehat{\chi})^{T}$ as long as it exists. When $b\geqslant 0$ we have already seen that $J_{v,b}^{\omega}$ is strictly convex along any $C^{1,1}$ geodesics in $\mathcal{H}(X,\widehat{\chi})^{T}$ (cf. Lemma 2.16). On the other hand it is not the case with $b<0$ so that we should take $\varepsilon$ sufficiently small according to the initial data in order to obtain the regularity of the flow. In what follows we study the long-time behavior of (3.1) in two cases $b\geqslant 0$ and $b=-\varepsilon<0$ separately.
3.1. Long-time existence for $b=-\varepsilon<0$
Assume $b=-\varepsilon<0$. Then by (1.3) and (1.7) the constant $c=c_{\varepsilon}$ is given by
$c_{\varepsilon}=\frac{n\alpha\cdot\beta^{n-1}-\varepsilon\alpha^{n}}{\beta^{n}}=c_{X}-\varepsilon\frac{\alpha^{n}}{\beta^{n}}.$ (3.2)
Fix $\phi_{0}\in\mathcal{H}(X,\widehat{\chi})^{T}$ and take a constant $K_{0}>0$ so that $Q_{\omega}(\chi_{0})=Q(A_{0})<K_{0}$. We set $\varepsilon_{5}:=\min\{\varepsilon_{1}(n,K_{0}+3C_{\theta}),\varepsilon_{2}(n,K_{0}+3C_{\theta}),\varepsilon_{3}(n,K_{0},C_{\theta})\}$ and assume that $\varepsilon\in(0,\varepsilon_{5})$, where the constants $\varepsilon_{1}$, $\varepsilon_{2}$, $\varepsilon_{3}$, $C_{\theta}$ are given in Proposition 2.3, Proposition 2.4, Lemma 2.5 and (2.2) respectively. Then the flow
$\frac{d}{dt}\phi_{t}=-F_{-\varepsilon}(A_{t})+c_{\varepsilon}+\theta_{v}^{X}(\chi_{t})$ (3.3)
is parabolic at $t=0$ from the choice of $\varepsilon$, so the short-time existence follows from standard theory. Let $T_{\max}$ be the maximal existence time of $\phi_{t}\in\mathcal{H}(X,\widehat{\chi})^{T}$. We introduce the operator $\square_{-\varepsilon,t}$ by
$\square_{-\varepsilon,t}:=\frac{d}{dt}+\partial_{ij}F_{-\varepsilon}(A_{t})g^{i\bar{k}}\nabla_{j}\nabla_{\bar{k}}-\mathrm{Re}(v),$
where $\nabla$ denotes the Chern connection with respect to $\omega$.
Lemma 3.1.
Along the flow $\phi_{t}$ for $t\in[0,T_{\max})$ we have the following:
  1. (1)
    $|\frac{d}{dt}\phi_{t}|<C$.
  2. (2)
    $|\phi_{t}|\leqslant C(t+1)$.
  3. (3)
    $Q(A_{t})<K_{0}+3C_{\theta}$.
Proof.
Differentiating the equation (3.3) we have
$\square_{-\varepsilon,t}\frac{d}{dt}\phi_{t}=0$
for all $t\in[0,T_{\max})$. So the maximum principle shows that
$\inf_{X}\frac{d}{dt}\phi_{0}\leqslant\frac{d}{dt}\phi_{t}\leqslant\sup_{X}\frac{d}{dt}\phi_{0}$
for all $t\in[0,T_{\max})$ which shows (1), and (2) by integrating in $t$. Moreover, the lower bound of $\frac{d}{dt}\phi_{t}$ implies that
$F_{-\varepsilon}(A_{t})\leqslant\theta_{v}^{X}(\chi_{t})+\sup_{X}\big(F_{-\varepsilon}(A_{0})-\theta_{v}^{X}(\chi_{0})\big)<K_{0}+2C_{\theta}.$
Then it follows from Lemma 2.5 that $Q(A_{t})<K_{0}+3C_{\theta}$ for all $t\in[0,T_{\max})$, and hence we have (3). ∎
Lemma 3.2.
The modified $J$-flow admits a solution $\phi_{t}$ for all $t\in[0,\infty)$.
Proof.
It follows from Lemma 3.1 that the function $F_{-\varepsilon}$ is convex as long as the flow exists. Therefore, by the Evans–Krylov estimate [Kry76, Kry82], the Schauder estimate, and the standard bootstrapping argument, the $C^{\infty}$-estimate for $\phi_{t}$ reduces to a bound on $|\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{t}|$. We take normal coordinates of $\omega$ at a point so that $\chi_{t}$ is diagonal with eigenvalues $\lambda_{1},\ldots,\lambda_{n}$. By using (2.1) we compute
$\nabla_{i}\nabla_{\bar{i}}F_{-\varepsilon}(A_{t})$ $=\nabla_{\bar{i}}\big(\partial_{pq}F_{-\varepsilon}(A_{t})\cdot\nabla_{i}A_{q}^{p}\big)$
$=\partial_{pq}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}\nabla_{i}A_{q}^{p}+\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}A_{s}^{r}\cdot\nabla_{i}A_{q}^{p}$
$=\partial_{pq}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}\nabla_{i}h_{q\bar{p}}+\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}h_{s\bar{r}}\cdot\nabla_{i}h_{q\bar{p}}$
$=\partial_{pp}F_{-\varepsilon}(A_{t})\cdot(\nabla_{p}\nabla_{\bar{p}}h_{i\bar{i}}+\operatorname{Rm}\ast h)+\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}h_{s\bar{r}}\cdot\nabla_{i}h_{q\bar{p}},$
where $\operatorname{Rm}$ denotes the Riemannian curvature of $\omega$. By the definition of Hamiltonians $\theta_{v}^{X}(\chi_{t})$ we have
$\nabla_{i}\nabla_{\bar{i}}\theta_{v}^{X}(\chi_{t})$ $=\nabla_{i}(v^{k}h_{k\bar{i}})$
$=\nabla_{i}v^{k}\cdot h_{k\bar{i}}+v^{k}\nabla_{i}h_{k\bar{i}}$
$=\nabla_{i}v^{k}\cdot h_{k\bar{i}}+v^{k}\nabla_{i}\widehat{h}_{k\bar{i}}+v^{k}\nabla_{i}\phi_{k\bar{i}}$
$=\nabla_{i}v^{k}\cdot h_{k\bar{i}}+v^{k}\nabla_{i}\widehat{h}_{k\bar{i}}+v^{k}\nabla_{k}\phi_{i\bar{i}}$
$\leqslant C_{1}|h|+v(\Delta_{g}\phi)+C_{1},$
where $\widehat{h}$ denotes the Riemannian metric corresponding to $\widehat{\chi}$, and we used $h_{k\bar{i}}=\widehat{h}_{k\bar{i}}+\phi_{k\bar{i}}$. By (3.3) we obtain
$\frac{d}{dt}\log\operatorname{Tr}_{\omega}\chi_{t}$
$=\frac{1}{\operatorname{Tr}_{\omega}\chi_{t}}\nabla_{i}\nabla_{\bar{i}}\big(-F_{-\varepsilon}(A_{t})+\theta_{v}^{X}(\chi_{t})\big)$
$=\frac{1}{\operatorname{Tr}_{\omega}\chi_{t}}\big(-\partial_{pp}F_{-\varepsilon}(A_{t})\cdot\nabla_{p}\nabla_{\bar{p}}h_{i\bar{i}}-\partial_{pp}F_{-\varepsilon}(A_{t})\cdot\operatorname{Rm}\ast h-\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}h_{s\bar{r}}\cdot\nabla_{i}h_{q\bar{p}}$
$+\nabla_{i}\nabla_{\bar{i}}\theta_{v}^{X}(\chi_{t})\big)$
$\leqslant\frac{1}{\operatorname{Tr}_{\omega}\chi_{t}}\big(-\partial_{pp}F_{-\varepsilon}(A_{t})\cdot\nabla_{p}\nabla_{\bar{p}}h_{i\bar{i}}+|\partial_{pp}F_{-\varepsilon}(A_{t})||\operatorname{Rm}||h|-\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}h_{s\bar{r}}\cdot\nabla_{i}h_{q\bar{p}}$
$+C_{1}|h|+v(\Delta_{g}\phi)+C_{1}\big).$
Also a direct computation shows that
$\nabla_{p}\nabla_{\bar{p}}\log\operatorname{Tr}_{\omega}\chi_{t}=\frac{\nabla_{p}\nabla_{\bar{p}}\operatorname{Tr}_{\omega}\chi_{t}}{\operatorname{Tr}_{\omega}\chi_{t}}-\frac{|\nabla_{p}\operatorname{Tr}_{\omega}\chi_{t}|^{2}}{(\operatorname{Tr}_{\omega}\chi_{t})^{2}}=\frac{\nabla_{p}\nabla_{\bar{p}}h_{i\bar{i}}}{\operatorname{Tr}_{\omega}\chi_{t}}-\frac{|\nabla_{i}h_{p\bar{i}}|^{2}}{(\operatorname{Tr}_{\omega}\chi_{t})^{2}},$
$v(\log\operatorname{Tr}_{\omega}\chi_{t})=\frac{v(\operatorname{Tr}_{\omega}\chi_{t})}{\operatorname{Tr}_{\omega}\chi_{t}}=\frac{1}{\operatorname{Tr}_{\omega}\chi_{t}}\big(v(\operatorname{Tr}_{\omega}\widehat{\chi})+v(\Delta_{g}\phi)\big)\geqslant\frac{1}{\operatorname{Tr}_{\omega}\chi_{t}}\big(v(\Delta_{g}\phi)-C_{2}\big).$
By using the strong convexity (Proposition 2.4) we observe that
$\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}h_{s\bar{r}}\cdot\nabla_{i}h_{q\bar{p}}+\partial_{pp}F_{-\varepsilon}(A_{t})\frac{|\nabla_{i}h_{p\bar{i}}|^{2}}{\operatorname{Tr}_{\omega}\chi_{t}}$
$\geqslant\partial_{pq}\partial_{rs}F_{-\varepsilon}(A_{t})\cdot\nabla_{\bar{i}}h_{s\bar{r}}\cdot\nabla_{i}h_{q\bar{p}}+\partial_{pp}F_{-\varepsilon}(A_{t})\frac{|\nabla_{i}h_{p\bar{j}}|^{2}}{\lambda_{j}}$
$\geqslant 0.$
Combining all things together we have
$\square_{-\varepsilon,t}\log\operatorname{Tr}_{\omega}\chi_{t}\leqslant\frac{1}{\operatorname{Tr}_{\omega}\chi_{t}}\big(|\partial_{pp}F_{-\varepsilon}(A_{t})||\operatorname{Rm}||h|+C_{1}|h|+C_{3}\big),$
where $\operatorname{Tr}_{\omega}\chi_{t}=\sum_{i}\lambda_{i}$, $|h|=(\sum_{i}\lambda_{i}^{2})^{1/2}$ and
$0<-\partial_{pp}F_{-\varepsilon}(A_{t})=-\partial_{p}f_{-\varepsilon}(\lambda)=\frac{1}{\lambda_{p}^{2}}-\frac{\varepsilon}{\lambda_{p}S_{n}(\lambda)}.$
Since $\lambda_{i}$ is uniformly bounded from blow this shows that the right hand side is bounded from above by some uniform constant $C_{4}>0$. Therefore if we set $G:=\log\operatorname{Tr}_{\omega}\chi_{t}-2C_{4}t$ then for any $T\in(0,T_{\max})$ and maximum point $(x_{0},t_{0})$ of $G$ on $X\times[0,T]$ with $t_{0}>0$ we must have
$0\leqslant\square_{-\varepsilon,t}G\leqslant-C_{4},$
but this is impossible. It follows that
$\operatorname{Tr}_{\omega}\chi_{t}\leqslant e^{2C_{4}t+C_{5}}$
on $0,T_{\max})$ which gives the uniform upper bound for $\lambda_{i}$. ∎
3.2. Convergence for $b\geqslant 0$
In this subsection we will give a proof of Theorem 1.1 by showing that the modified $J$-flow
$\frac{d}{dt}\phi_{t}=-F_{b}(A_{t})+c+\theta_{v}^{X}(\chi_{t})$ (3.4)
converges smoothly to the solution (1.6), under the assumption that there exists a subsolution $\widehat{\chi}\in\beta$ satisfying
[…]
which is chosen as a reference Kähler form. Define the operator
$\square_{b,t}:=\frac{d}{dt}+\partial_{ij}F_{b}(A_{t})g^{i\bar{k}}\nabla_{j}\nabla_{\bar{k}}-\mathrm{Re}(v),$
and let $T_{\max}$ be the maximal existence time of the flow.
Lemma 3.3.
There exists a uniform constant $C>0$ such that
$\bigg|\frac{d}{dt}\phi_{t}\bigg|+Q(A_{t})<C.$
In particular, the eigenvalues of $A_{t}$ have a uniform positive lower bound.
Proof.
The proof is similar to that of Lemma 3.1. ∎
Lemma 3.4.
There exists a uniform constant $N>0$, $C>0$ such that
$\operatorname{Tr}_{\omega}\chi_{t}\leqslant Ce^{N(\phi_{t}-\inf_{M\times[0,T]}\phi_{t})}$ (3.6)
for all $t\in[0,T]$ and $T\in[0,T_{\max})$.
Proof.
From the assumption we can choose a constant $\epsilon>0$ sufficiently small so that
$P_{\omega}(\widehat{\chi})<c+\theta_{v}^{X}(\widehat{\chi})-2\epsilon.$
We choose normal coordinates for $\omega$ so that $\chi_{t}$ is diagonal with entries $\lambda_{1}\leqslant\ldots\leqslant\lambda_{n}$. The metric $\widehat{\chi}$ may not be diagonal, but we denote its eigenvalues by $\mu_{1},\ldots,\mu_{n}$. As for the first part, the calculation in Lemma 3.2 can be applied without modification (indeed, one can easily check that the constants $C_{1},\ldots,C_{4}$ in the proof of Lemma 3.2 can all be chosen independently of $t$). Consequently, we obtain
$\square_{b,t}\log\operatorname{Tr}_{\omega}\chi_{t}\leqslant C_{1}.$
The following discussion is restricted to the case $b\geqslant 0$, and our goal is to obtain a $t$-independent upper bound for the eigenvalues $\lambda_{i}$. We first compute
$\square_{b,t}\phi_{t}=-F_{b}(A_{t})+c+\theta_{v}^{X}(\widehat{\chi})+\partial_{ii}F_{b}(A_{t})(h_{ii}-\widehat{h}_{i\bar{i}}).$
Hence if we set $G:=\log\operatorname{Tr}_{\omega}\chi_{t}-N\phi_{t}$ then
$\square_{b,t}G\leqslant N\epsilon+NF_{b}(A_{t})-Nc-N\theta_{v}^{X}(\widehat{\chi})+N\partial_{ii}F_{b}(A_{t})(\widehat{h}_{i\bar{i}}-h_{i\bar{i}}),$
where the large constant $N>0$ is taken so that $C_{1}\leqslant N\epsilon$. For any fixed $T\in(0,T_{\max})$ the maximum principle shows that
$\epsilon\leqslant F_{b}(A_{t})-c-\theta_{v}^{X}(\widehat{\chi})+\partial_{ii}F_{b}(A_{t})(\widehat{h}_{i\bar{i}}-h_{i\bar{i}})$
at the maximum point $(x_{0},t_{0})$ of $G$ on $X\times[0,T]$ if $t_{0}>0$. In particular,
$P_{\omega}(\widehat{\chi})+\epsilon\leqslant F_{b}(A_{t})+\partial_{ii}F_{b}(A_{t})(\widehat{h}_{i\bar{i}}-h_{i\bar{i}})$
from our choice of $\epsilon$, which can be written as
$\max_{k}\sum_{i\neq k}\frac{1}{\mu_{i}}+\epsilon\leqslant f_{b}(\lambda)+\partial_{i}f_{b}(\lambda)\cdot(\widehat{h}_{i\bar{i}}-\lambda_{i}).$
Following [CS17, Section 2] we define the function
$\widetilde{f}_{b}(\lambda_{1},\ldots,\lambda_{n-1}):=\lim_{\lambda_{n}\to\infty}f_{b}(\lambda_{1},\ldots,\lambda_{n})=\sum_{i=1}^{n-1}\frac{1}{\lambda_{i}}.$
By using the uniform positive lower bound for $\lambda_{i}$, one can easily observe that for any constant $\tau>0$ there exists $K=K(\tau)>0$ such that if $\lambda_{n}\geqslant K$ then
$f_{b}(\lambda_{1},\ldots,\lambda_{n})\leqslant\widetilde{f}_{b}(\lambda_{1},\ldots,\lambda_{n-1})+\tau,$
$\sum_{i=1}^{n-1}\big(\partial_{i}f_{b}(\lambda_{1},\ldots,\lambda_{n})-\partial_{i}\widetilde{f}_{b}(\lambda_{1},\ldots,\lambda_{n-1})\big)\cdot(\widehat{h}_{i\bar{i}}-\lambda_{i})=\sum_{i=1}^{n-1}\frac{b}{\lambda_{1}\cdots\lambda_{n}}\bigg(1-\frac{\widehat{h}_{i\bar{i}}}{\lambda_{i}}\bigg)\leqslant\tau,$
$\partial_{n}f_{b}(\lambda)\cdot(\widehat{h}_{n\bar{n}}-\lambda_{n})=\frac{1}{\lambda_{n}}\bigg(1+\frac{b}{\lambda_{1}\cdots\lambda_{n-1}}\bigg)\bigg(1-\frac{\widehat{h}_{n\bar{n}}}{\lambda_{n}}\bigg)\leqslant\tau.$
Combining with the convexity of $\widetilde{f}_{b}$ yields that
$f_{b}(\lambda)+\partial_{i}f_{b}(\lambda)\cdot(\widehat{h}_{i\bar{i}}-\lambda_{i})$
$\leqslant\widetilde{f}_{b}(\lambda_{1},\ldots,\lambda_{n-1})+\sum_{i=1}^{n-1}\partial_{i}\widetilde{f}_{b}(\lambda_{1},\ldots,\lambda_{n-1})\cdot(\widehat{h}_{i\bar{i}}-\lambda_{i})+3\tau$
$\leqslant\widetilde{f}_{b}(\widehat{h}_{1\bar{1}},\ldots,\widehat{h}_{n-1\overline{n-1}})+3\tau$
$\leqslant\max_{k}\sum_{i\neq k}\frac{1}{\mu_{i}}+3\tau,$
where in the last inequality we used the fact that diagonal entries $(\widehat{h}_{1\bar{1}},\ldots,\widehat{h}_{n\bar{n}})$ lies in the convex full of the permutations of the $\mu_{i}$ by the Schur–Horn theorem. Eventually if we choose $\tau=\epsilon/6$ then we have $\lambda_{n}\leqslant K(\tau)$, which yields the desired estimate. ∎
Once we obtain (3.6) then we can show the uniform $C^{0}$-estimate for $\phi_{t}$ by using exactly the same argument as [SW08, Wei04, Wei06] since it does not depend on the equation (3.4). The higher order estimate follows from the Evans–Krylov estimate [Kry76, Kry82], the Schauder estimate and the standard bootstrapping argument. Consequently, all derivatives of $\phi_{t}$ are uniformly bounded independent of $t$, and in particular, the flow $\phi_{t}$ exists for all $t\in[0,\infty)$. Now we establish the following:
Theorem 3.5.
Assume that the reference Kähler form $\widehat{\chi}\in\beta$ is a subsolution (3.5). Then along the modified $J$-flow (3.4), the Kähler form $\chi_{\phi_{t}}$ converges smoothly to a solution of (1.6) as $t\to\infty$.
Proof.
For $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ we define
$E_{v,b}^{\omega}(\phi):=\int_{X}\big(F_{\omega,b}(\chi_{\phi})-c-\theta_{v}^{X}(\chi_{\phi})\big)^{2}\chi_{\phi}^{n}.$
We will compute the derivative of $E_{v,b}^{\omega}$ along the flow. To simplify the notations we set
$\sigma:=\frac{d}{dt}\phi_{t}=-F_{b}(A_{t})+c+\theta_{v}^{X}(\chi_{t}),\quad\nu:=\frac{\omega^{n}}{\chi_{t}^{n}}.$
In normal coordinates of $\chi_{t}$ we compute
$\frac{d}{dt}\sigma=h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{j}}\sigma_{k\bar{\ell}}+b\nu\Delta_{t}\sigma+v(\sigma).$
Hence we have
$\frac{d}{dt}E_{v,b}^{\omega}$ $=2\int_{X}\sigma\frac{d}{dt}\sigma\chi_{t}^{n}+\int_{X}\sigma^{2}\Delta_{t}\sigma\chi_{t}^{n}$
$=2\int_{X}\bigg(\sigma\frac{d}{dt}\sigma-\sigma|\nabla\sigma|_{h}^{2}\bigg)\chi_{t}^{n}$
$=2\int_{X}\big(h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{j}}\sigma\sigma_{k\bar{\ell}}+b\nu h^{k\bar{\ell}}\sigma\sigma_{k\bar{\ell}}+\sigma v(\sigma)-h^{k\bar{\ell}}\sigma\sigma_{k}\sigma_{\bar{\ell}}\big)\chi_{t}^{n}$
$=-2\int_{X}\big(h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{j}}\sigma_{k}\sigma_{\bar{\ell}}+h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{j},\bar{\ell}}\sigma\sigma_{k}+b\nu_{\bar{\ell}}h^{k\bar{\ell}}\sigma\sigma_{k}+b\nu h^{k\bar{\ell}}\sigma_{k}\sigma_{\bar{\ell}}-\sigma v(\sigma)+h^{k\bar{\ell}}\sigma\sigma_{k}\sigma_{\bar{\ell}}\big)\chi_{t}^{n}$
$=-2\int_{X}\big(h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{j}}\sigma_{k}\sigma_{\bar{\ell}}+h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{\ell},\bar{j}}\sigma\sigma_{k}+b\nu_{\bar{\ell}}h^{k\bar{\ell}}\sigma\sigma_{k}+b\nu h^{k\bar{\ell}}\sigma_{k}\sigma_{\bar{\ell}}-\sigma v(\sigma)+h^{k\bar{\ell}}\sigma\sigma_{k}\sigma_{\bar{\ell}}\big)\chi_{t}^{n}$
$=-2\int_{X}\big(h^{i\bar{\ell}}h^{k\bar{j}}g_{i\bar{j}}\sigma_{k}\sigma_{\bar{\ell}}+h^{k\bar{j}}(\operatorname{Tr}_{\chi_{t}}\omega)_{\bar{j}}\sigma\sigma_{k}+b\nu_{\bar{\ell}}h^{k\bar{\ell}}\sigma\sigma_{k}+b\nu h^{k\bar{\ell}}\sigma_{k}\sigma_{\bar{\ell}}-\sigma v(\sigma)$
$-h^{k\bar{\ell}}\sigma\sigma_{k}(\operatorname{Tr}_{\chi_{t}}\omega+b\nu-\theta_{v}^{X}(\chi_{t}))_{\bar{\ell}}\big)\chi_{t}^{n}$
$=-2\bigg(\int_{X}|\nabla\sigma|_{\eta}^{2}\chi_{t}^{n}+b\int_{X}|\nabla\sigma|_{h}^{2}\omega^{n}\bigg)$
$\leqslant 0,$
where the Hermitian metric $\eta$ is defined by $\eta_{i\bar{j}}:=g^{k\bar{\ell}}h_{i\bar{\ell}}h_{k\bar{j}}$. Since we have already obtained a uniform $C^{\infty}$-estimate for $\phi_{t}$, we know that
$\int_{0}^{\infty}\bigg(\int_{X}|\nabla\sigma|_{\eta}^{2}\chi_{t}^{n}+b\int_{X}|\nabla\sigma|_{h}^{2}\omega^{n}\bigg)dt<+\infty.$
By passing to a subsequence $t_{k}$ we get
$\int_{X}|\nabla\sigma(t_{k})|_{\eta}^{2}\chi_{t_{k}}^{n}+b\int_{X}|\nabla\sigma(t_{k})|_{h}^{2}\omega^{n}\to 0,$
and hence $\sigma(t_{k})\to 0$ by $\int_{X}\sigma(t_{k})\chi_{t_{k}}^{n}=0$. This shows that the limit $\chi_{t_{\infty}}$ satisfies the generalized equation
$\operatorname{Tr}_{\chi_{t_{\infty}}}\omega+b\frac{\omega^{n}}{\chi_{t_{\infty}}^{n}}=c+\theta_{v}^{X}(\chi_{t_{\infty}}),$
where the Kähler form $\chi_{t_{\infty}}$ does not depend on the choice of subsequences, thanks to the strict convexity property of $J_{v,b}^{\omega}$ (cf. Proposition 2.16). Thus we conclude that $\chi_{\phi_{t}}$ converges smoothly to the unique solution of (1.6). ∎
In particular, Theorem 3.5 provides a proof of Theorem 1.1.
4. Coercivity of the modified $J$-functional
4.1. Coercivity implies existence
In this section we will prove Theorem 1.2. As in [CS17, Proposition 21] a first crucial observation is that if $J_{v}^{\omega}$ is coercive then so is $J_{v,-\varepsilon}^{\omega}$ for sufficiently small $\varepsilon>0$. For a fixed $\omega\in\alpha$ we take a reference Kähler form $\widehat{\chi}\in\beta$ as a solution to the complex Monge–Ampère equation
$\widehat{\chi}^{n}=\frac{\beta^{n}}{\alpha^{n}}\omega^{n},$
which is known to exist by Yau’s theorem [Yau78].
Proposition 4.1.
Suppose that there exist uniform constants $\delta_{c}>0$, $B_{c}>0$ such that
$J_{v}^{\omega}(\phi)\geqslant\delta_{c}I(\phi)-B_{c}$
for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$. Then there exists $\varepsilon_{6}=\varepsilon_{6}(n,\alpha,\beta,\delta_{c})>0$ such that for all $\varepsilon\in(0,\varepsilon_{6})$ and $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ we have
$J_{v,-\varepsilon}^{\omega}(\phi)\geqslant\frac{\delta_{c}}{2}I(\phi)-B_{c}.$
Proof.
If we subtract $J_{v}^{\omega}$ from $J_{v,-\varepsilon}^{\omega}$ the integrands involving Hamiltonians cancel each other out. We can therefore proceed as in [CS17, Proposition 21] to see that
$J_{v,-\varepsilon}^{\omega}(\phi)$ $=J_{v}^{\omega}(\phi)+\varepsilon\int_{0}^{1}\int_{X}\phi\bigg(\frac{\alpha^{n}}{\beta^{n}}\chi_{t\phi}^{n}-\omega^{n}\bigg)dt$
$=J_{v}^{\omega}(\phi)+\varepsilon\frac{\alpha^{n}}{\beta^{n}}\int_{0}^{1}\int_{X}\phi(\chi_{t\phi}^{n}-\widehat{\chi}^{n})dt$
$=J_{v}^{\omega}(\phi)+\varepsilon\frac{\alpha^{n}}{\beta^{n}}\int_{0}^{1}\int_{X}t\phi\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi\wedge(\chi_{t\phi}^{n-1}+\chi_{t\phi}^{n-2}\wedge\widehat{\chi}+\cdots+\widehat{\chi}^{n-1})dt$
$\geqslant J_{v}^{\omega}(\phi)-\varepsilon\frac{\sqrt{-1}}{2\pi}\frac{\alpha^{n}}{\beta^{n}}\int_{0}^{1}\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge(\chi_{t\phi}^{n-1}+\chi_{t\phi}^{n-2}\wedge\widehat{\chi}+\cdots+\widehat{\chi}^{n-1})dt.$
By using the coercivity estimate $J_{v}^{\omega}(\phi)\geqslant\delta_{c}I(\phi)-B_{c}$ we have
$J_{v,-\varepsilon}^{\omega}(\phi)$ $\geqslant\delta_{c}\frac{\sqrt{-1}}{2\pi}\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge(\chi_{\phi}^{n-1}+\chi_{\phi}^{n-2}\wedge\widehat{\chi}+\cdots+\chi^{n-1})$ (4.1)
$-\varepsilon\frac{\sqrt{-1}}{2\pi}\frac{\alpha^{n}}{\beta^{n}}\int_{0}^{1}\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge(\chi_{t\phi}^{n-1}+\chi_{t\phi}^{n-2}\wedge\widehat{\chi}+\cdots+\widehat{\chi}^{n-1})dt-B_{c}.$
Since $\chi_{t\phi}=(1-t)\chi+t\chi_{\phi}$ the integrands involving $\chi_{t\phi}$ can be expressed as
$\chi_{t\phi}^{n-1-k}\wedge\widehat{\chi}^{k}=\sum_{i=0}^{n-1-k}p_{i}(t)\chi_{\phi}^{n-1-k-i}\wedge\widehat{\chi}^{k+i}$
for all $k=0,1,\ldots,n-1$, where $p_{i}(t):=\binom{n-1-k}{i}(1-t)^{i}t^{n-1-k-i}$. This shows that the second term in the right hand side of (4.1) can be absorbed in the first term if $\varepsilon$ is sufficiently small. So there exists $\varepsilon_{6}>0$ such that if $\varepsilon\in(0,\varepsilon_{6})$ then
$J_{v,-\varepsilon}^{\omega}(\phi)\geqslant\frac{\delta_{c}}{2}I(\phi)-B_{c}$
for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$. ∎
From now on, we assume that $J_{v}^{\omega}(\phi)\geqslant\delta_{c}I(\phi)-B_{c}$ holds for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$. Take a large number $K_{0}>0$ so that $Q_{\omega}(\widehat{\chi})<K_{0}$ and set $\varepsilon_{7}:=\min\{\varepsilon_{4}(n,K_{0}+3C_{\theta}),\varepsilon_{5},\varepsilon_{6}(n,\alpha,\beta,\delta_{c})\}$, where $\varepsilon_{4}$ is defined in Lemma 2.7 and $\varepsilon_{5}$ is defined in Section 3.1. Then by Lemma 3.1 and Lemma 3.2 the modified $J$-flow starting from $0\in\mathcal{H}(X,\widehat{\chi})^{T}$ exists for all $t\in[0,\infty)$ with a uniform bound
$Q_{\omega}(\chi_{\phi_{t}})<K_{0}+3C_{\theta}.$ (4.2)
By the coercivity estimate, there is a sequence $t_{k}\to\infty$ such that
$\lim_{k\to\infty}\frac{d}{dt}J_{v,-\varepsilon}^{\omega}(\phi_{t_{k}})=0,$
and hence
$\lim_{k\to\infty}\int_{X}\big(F_{\omega,-\varepsilon}(\chi_{\phi_{t_{k}}})-c_{\varepsilon}-\theta_{v}^{X}(\chi_{\phi_{t_{k}}})\big)^{2}\omega^{n}=0,$
where we used the uniform lower bound for eigenvalues along the flow $\phi_{t}$. We normalize $\phi_{t_{k}}$ so that
$u_{k}:=\phi_{t_{k}}-\sup_{X}\phi_{t_{k}}.$
Then
$\lim_{k\to\infty}\int_{X}\big(F_{\omega,-\varepsilon}(\chi_{u_{k}})-c_{\varepsilon}-\theta_{v}^{X}(\chi_{u_{k}})\big)^{2}\omega^{n}=0.$
[…]
$I(u_{k})\leqslant\frac{J_{v,-\varepsilon}^{\omega}(u_{k})+B_{c}}{\delta_{c}/2}\leqslant\frac{J_{v,-\varepsilon}^{\omega}(0)+B_{c}}{\delta_{c}/2}\leqslant C$ (4.3)
[…]
4.1.1. Local smoothing
The Kähler current $\chi_{u_{\infty}}$ can be regarded as a solution to (1.2) in a weak sense. Now we establish estimates for local smoothing of $\chi_{u_{\infty}}$ in Euclidean balls. The main difference from the usual $J$-equation lies in the smoothing estimates for the Hamiltonian, which are not straightforward. Even if $v$ is locally approximated by a constant-coefficient vector field, controlling the resulting error term requires a uniform gradient bound $|\nabla u_{k}|_{g}\leqslant C$. While this error term could be eliminated using the rectification theorem, this approach can not be applied directly since $v$ must have a zero on $X$. To address this issue, one may consider a product manifold $\mathcal{X}:=X\times\mathbb{C}$ instead of $X$: let $\omega_{\mathbb{C}}:=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}|\tau|^{2}$ be the standard Kähler form on $\mathbb{C}$ and $w:=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$ a holomorphic vector field on $\mathbb{C}$ whose real part generates the action of the real torus $T^{\prime}\simeq\mathbb{S}^{1}$ on $\mathbb{C}$ by rotations. Set
$\omega_{\mathcal{X}}:=\omega+\omega_{\mathbb{C}},\quad\widehat{\chi}_{\mathcal{X}}:=\widehat{\chi}+\varepsilon^{-2}\omega_{\mathbb{C}},$
where each term on the right-hand side is the pullback, via the projection, of a Kähler form defined on each component (we will also omit writing the pullback via the projection in later arguments). Similarly, we extend $v$ and $w$ to holomorphic vector fields on $\mathcal{X}$ by means of the product structure, and obtain the product torus action $T\times T^{\prime}$ on $\mathcal{X}$ as well. Define
$\mathcal{X}_{k}:=X\times\{\tau\in\mathbb{C}|2^{-k}<|\tau|<2^{k}\},\quad k=1,2,3.$
Since the holomorphic vector field $v+w$ has no zeros on $\mathcal{X}_{3}$, for any fixed $p\in\mathcal{X}_{3}$ we can choose local homomorphic coordinates $(w^{1},\ldots,w^{n+1})$ centered at $p$ such that $\frac{\partial}{\partial w^{1}}=v+w$ by the rectification theorem. In this coordinates, after adjusting the affine term, the local potential function $\phi$ of $\omega_{\mathcal{X}}$ is expressed as $\phi(w)=\overline{w}^{T}Gw+O(|w|^{3})$ for some positive definite Hermitian matrix $G=(G_{i\bar{j}})$. In particular, we have $|v+w|_{g_{\mathcal{X}}}^{2}(p)=G_{1\bar{1}}$, where $g_{\mathcal{X}}$ denotes the Riemannian metric with respect to $\omega_{\mathcal{X}}$. Take $B:=G^{1/2}$ so that the local potential function $\phi$ in the coordinates $z=Bw$ becomes $\phi(z)=|z|^{2}+O(|z|^{3})$. The holomorphic vector field $v+w$ still has constant coefficients in $(z^{1},\ldots,z^{n+1})$, which we denote by $c(v+w)\in\mathbb{C}^{n+1}$. Also a simple computation shows that $|c(v+w)|=|v+w|_{g_{\mathcal{X}}}(p)\leqslant\sup_{\mathcal{X}_{3}}|v+w|_{\omega_{\mathcal{X}}}$.
Taking this into the account, we can choose sufficiently small constants $R>0$, $\sigma>0$ and a finite covering $\{B_{j,4R}\}_{j\in\mathcal{J}}$ of $\mathcal{X}_{2}$ satisfying the following properties:
  • Each $B_{j,4R}$ is a coordinate chart of $\mathcal{X}$ biholomorphic to a Euclidean ball $B_{4R}(0)$ of radius $4R$ centered at the origin in $\mathbb{C}^{n+1}$, and $\bigcup_{j\in\mathcal{J}}B_{j,4R}\subset\mathcal{X}_{3}$.
  • The family $\{B_{j,R}\}_{j\in\mathcal{J}}$ also forms a covering of $\mathcal{X}_{2}$, where each $B_{j,R}$ is biholomorphic to the Euclidean ball $B_{R}(0)\subset\mathbb{C}^{n+1}$.
  • On each $B_{j,4R}$, we take local potential functions $\phi_{\omega,j}$, $\phi_{\widehat{\chi},j}$ as
    $\omega_{\mathcal{X}}=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{\omega,j},\quad\widehat{\chi}_{\mathcal{X}}=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{\widehat{\chi},j}.$
    We may assume that in the coordinates $(z^{1},\ldots,z^{n+1})$ of $B_{j,4R}$, the holomorphic vector field $v+w$ has constant coefficients $c(v+w)\in\mathbb{C}^{n+1}$ and $\phi_{\omega,j}(z)=|z|^{2}+O(|z|^{3})$. In particular, since $\partial\phi_{\omega,j}(z)=O(|z|)$ and $\partial\bar{\partial}O(|z|^{3})=O(|z|)$, we may further assume that
    $(1-\sigma)\omega_{j}\leqslant\omega_{\mathcal{X}}\leqslant(1+\sigma)\omega_{j},$
    $\big|\phi_{\omega,j}-|z|^{2}\big|\leqslant\frac{1}{10000}R^{2},\quad|c(v+w)|+|\nabla\phi_{\omega,j}|\leqslant C$
    for some constant $C>0$ depending only on $v$, $w$ and $\omega_{\mathcal{X}}$, where $\omega_{j}$ denotes the Kähler form on $B_{j,4R}$ with constant coefficients corresponding to the Euclidean metric $g_{j}$. Also we normalize $\phi_{\widehat{\chi},j}$ so that $\phi_{\widehat{\chi},j}(0)=0$.
For each $j\in\mathcal{J}$ and any locally $L^{1}$-integrable function $f$ on $B_{j,4R}$, the smoothing of $f$ at scale $r\in(0,R)$ is defined by
$f^{(r)}(z):=\int_{\mathbb{C}^{n+1}}r^{-2n-2}\rho\bigg(\frac{|y|}{r}\bigg)f(z-y)d\operatorname{Vol}_{g_{j}}(y),\quad z\in B_{j,3R},$ (4.4)
where $\rho(t)$ is a smooth, non-negative function supported on $[0,1]$, constant on $[0,1/2]$, and normalized so that
$\int_{B_{1}(0)}\rho(|y|)d\operatorname{Vol}_{g_{j}}(y)=1.$
Furthermore, if $f$ is smooth, then $f^{(r)}\to f$ in $C^{\infty}$ as $r\to 0$. So once $\varepsilon$, $R$, $\sigma$ and $\{B_{j,4R}\}_{j\in\mathcal{J}}$ are fixed, there exists a constant $r_{0}\in(0,R)$ such that for all $r\in(0,r_{0})$ and $j\in\mathcal{J}$ we have
$\big|\mathrm{Re}(v+w)(\phi_{\widehat{\chi},j}^{(r)}-\phi_{\widehat{\chi},j})\big|\leqslant\sigma,\quad\big|\theta_{v}^{X}(\widehat{\chi})^{(r)}-\theta_{v}^{X}(\widehat{\chi})\big|\leqslant\sigma$ (4.5)
on $B_{j,3R}$. For each $k$ and $j\in\mathcal{J}$, define
$u_{k,j}:=\phi_{\widehat{\chi},j}+u_{k},\quad u_{\infty,j}:=\phi_{\widehat{\chi},j}+u_{\infty}.$
By Fubini’s theorem, we then have $u_{k,j}\to u_{\infty,j}$ in $L_{\rm loc}^{1}(B_{j,4R})$ as $k\to\infty$. Also, we have
$\widehat{\chi}_{\mathcal{X}}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k}=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j},\quad\widehat{\chi}_{\mathcal{X}}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty}=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j},$
where the second equality holds in the sense of currents.
The following lemma is crucial to our argument. The key is to estimate $R_{\omega_{\mathcal{X}}}$ rather than $P_{\omega_{\mathcal{X}}}$. Indeed, a simple computation shows that $P_{\omega_{\mathcal{X}}}$ produces an $O(\varepsilon)$ correction term, and hence is ineffective. However, in view of Proposition 2.2 (2), it suffices to estimate $R_{\omega_{\mathcal{X}}}$ to construct a subsolution on $X$.
Lemma 4.2.
For any $\varepsilon\in(0,\varepsilon_{7})$ we have
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\leqslant F_{\omega,-\varepsilon}(\chi_{u_{k}})+\varepsilon^{2},$
where the quantity in the left-hand side is calculated on $\mathcal{X}$.
Proof.
Let $\lambda_{1}\leqslant\ldots\leqslant\lambda_{n}$ be the eigenvalues of $\chi_{u_{k}}$ with respect to $\omega$. Then the eigenvalues of $\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}=\chi_{u_{k}}+\varepsilon^{-2}\omega_{\mathbb{C}}$ with respect to $\omega_{\mathcal{X}}$ are $\lambda_{1},\ldots,\lambda_{n},\varepsilon^{-2}$. If $\lambda_{n-1}\leqslant\varepsilon^{-2}$, then
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)=\sum_{i=1}^{n-1}\frac{1}{\lambda_{i}}=P_{\omega}(\chi_{u_{k}}).$
Otherwise we have
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)=\sum_{i=1}^{n-2}\frac{1}{\lambda_{i}}+\varepsilon^{2}\leqslant P_{\omega}(\chi_{u_{k}})+\varepsilon^{2}.$
Thus combining with Lemma 2.7 and (4.2) we obtain the desired estimate. ∎
Lemma 4.3.
There are constants $\varepsilon_{8}\in(0,\varepsilon_{7})$ and $\delta=\delta(K_{0},C_{\theta})>0$ such that the following holds: for any $\varepsilon\in(0,\varepsilon_{8})$ there exists constant $\sigma_{0}>0$ such that if $\sigma\in(0,\sigma_{0})$ then
  1. (1)
    $R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v+w)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j})\leqslant c_{X}-\delta\varepsilon$,
  2. (2)
    $Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)\leqslant K_{0}+5C_{\theta}$
on $B_{j,3R}$ for all $j\in\mathcal{J}$ and $r\in(0,r_{0})$, where $r_{0}$ is chosen in (4.5).
Proof.
In what follows, let $C$ denote a constant depending only on $K_{0}$ and $C_{\theta}$, which may vary from line to line. First, by (4.2) we compute
$Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)=Q_{\omega}(\chi_{u_{k}})+\varepsilon^{2}<K_{0}+3C_{\theta}+\varepsilon^{2}.$
Thus by Lemma 2.6 we know that
$\bigg|Q_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)-Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\bigg|\leqslant C\sigma,$ (4.6)
$\bigg|R_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)-R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\bigg|\leqslant C\sigma.$ (4.7)
By using the convexity of $Q_{\omega_{j}}$, (4.6) and the fact that the coefficients of $\omega_{j}$ are constants we have
$\frac{1}{1+\sigma}Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}^{(r)}\bigg)(z)$ $\leqslant Q_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}^{(r)}\bigg)(z)$
$\leqslant\int_{\mathbb{C}^{n+1}}r^{-2n-2}\rho\bigg(\frac{|y|}{r}\bigg)Q_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)(z-y)d\operatorname{Vol}_{g_{j}}(y)$
$\leqslant\int_{\mathbb{C}^{n+1}}r^{-2n-2}\rho\bigg(\frac{|y|}{r}\bigg)Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)(z-y)d\operatorname{Vol}_{g_{j}}(y)+C\sigma$
$\leqslant K_{0}+3C_{\theta}+\varepsilon^{2}+C\sigma.$
So if $\sigma<\varepsilon^{2}$ and $\varepsilon$ is small so that
$\varepsilon^{2}\leqslant\frac{2C_{\theta}}{K_{0}+3C_{\theta}+2C+2},$
then
$Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}^{(r)}\bigg)\leqslant K_{0}+5C_{\theta}.$
By letting $k\to\infty$ we have
$Q_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)\leqslant K_{0}+5C_{\theta},$
which shows (2). Next, since $\omega_{j}$ and $v+w$ have constant coefficients, by using (4.7), convexity of $R_{\omega_{j}}$, $\mathrm{Re}(w)(u_{k})=0$ and Lemma 4.2 we have
$R_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}^{(r)}\bigg)(z)-\mathrm{Re}(v+w)(u_{k}^{(r)})(z)$
$\leqslant\int_{\mathbb{C}^{n+1}}r^{-2n-2}\rho\bigg(\frac{|y|}{r}\bigg)\bigg[R_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)(z-y)-\mathrm{Re}(v+w)(u_{k})(z-y)\bigg]d\operatorname{Vol}_{g_{j}}(y)$
$\leqslant\int_{\mathbb{C}^{n+1}}r^{-2n-2}\rho\bigg(\frac{|y|}{r}\bigg)\bigg[R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)(z-y)-\mathrm{Re}(v)(u_{k})(z-y)\bigg]d\operatorname{Vol}_{g_{j}}(y)+C\sigma$
$\leqslant\int_{\mathbb{C}^{n+1}}r^{-2n-2}\rho\bigg(\frac{|y|}{r}\bigg)\big[F_{\omega,-\varepsilon}(\chi_{u_{k}})(z-y)-\mathrm{Re}(v)(u_{k})(z-y)\big]d\operatorname{Vol}_{g_{j}}(y)+C\sigma+\varepsilon^{2}$
[…]
$R_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)-\mathrm{Re}(v+w)(u_{\infty}^{(r)})$ $\leqslant c_{\varepsilon}+\big(\theta_{v}^{X}(\widehat{\chi})\big)^{(r)}+C\sigma+\varepsilon^{2}$ (4.8)
$\leqslant c_{\varepsilon}+\theta_{v}^{X}(\widehat{\chi})+C\sigma+\varepsilon^{2}$
on $B_{j,3R}$ for all $j\in\mathcal{J}$ and $r\in(0,r_{0})$. Meanwhile, by using (2) and Lemma 2.6, we find that
$\bigg|R_{\omega_{j}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)-R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)\bigg|\leqslant C\sigma.$ (4.9)
Combining (4.8) and (4.9), we obtain
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v+w)(u_{\infty}^{(r)})\leqslant c_{\varepsilon}+C\sigma+\varepsilon^{2}.$
By definition, we have $u_{\infty,j}^{(r)}=(\phi_{\widehat{\chi},j}+u_{\infty})^{(r)}=\phi_{\widehat{\chi},j}^{(r)}+u_{\infty}^{(r)}$, which implies that
$\mathrm{Re}(v+w)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j})=\mathrm{Re}(v+w)(u_{\infty}^{(r)})+\mathrm{Re}(v+w)(\phi_{\widehat{\chi},j}^{(r)}-\phi_{\widehat{\chi},j})\geqslant\mathrm{Re}(v+w)(u_{\infty}^{(r)})-\sigma$
from our choice of $r\in(0,r_{0})$. Consequently, we obtain
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j})\leqslant c_{\varepsilon}+C\sigma+\varepsilon^{2}.$
Recalling that $c_{\varepsilon}=c_{X}-\varepsilon\frac{\alpha^{n}}{\beta^{n}}$ by (3.2), we obtain (1) for some $\delta>0$ provided that $\sigma<\varepsilon^{2}$ and $\varepsilon$ is sufficiently small. ∎
4.1.2. Gluing argument
We now perform suitable modifications of the functions $u_{\infty,j}^{(r)}$ and glue them together to construct a smooth potential function on $\mathcal{X}_{1}$. To this end, as explained in Section 2.3, we need to verify the gluing conditions. Let $\lambda_{1}\leqslant\cdots\leqslant\lambda_{n+1}$ be the eigenvalues of $\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}$ with respect to $\omega_{\mathcal{X}}$. Then Lemma 4.3 (2) yields that
$\sum_{i=1}^{n+1}\frac{1}{\lambda_{i}}\leqslant K_{0}+5C_{\theta},$
which gives a uniform bound $\lambda_{i}\geqslant(K_{0}+5C_{\theta})^{-1}$. By using this we observe that the form $\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}-\varepsilon^{2}\omega_{\mathcal{X}}$ is Kähler and
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}-\varepsilon^{2}\omega_{\mathcal{X}}\bigg)$ $=\sum_{i=1}^{n-1}\frac{1}{\lambda_{i}-\varepsilon^{2}}$
$=\sum_{i=1}^{n-1}\frac{1}{\lambda_{i}}+\sum_{i=1}^{n-1}\frac{\varepsilon^{2}}{(\lambda_{i}-\varepsilon^{2})\lambda_{i}}$
$\leqslant\sum_{i=1}^{n-1}\frac{1}{\lambda_{i}}+C\varepsilon^{2}$
$=R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}\bigg)+C\varepsilon^{2}$
by decreasing $\varepsilon_{8}$ if necessary, where $C=C(K_{0},C_{\theta})>0$ is a constant. Also we have
$\mathrm{Re}(v+w)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j}-\varepsilon^{2}\phi_{\omega,j})$ $=\mathrm{Re}(v+w)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j})-\varepsilon^{2}\mathrm{Re}(v+w)(\phi_{\omega,j})$
$\geqslant\mathrm{Re}(v+w)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j})-C\varepsilon^{2},$
where we used $|c(v+w)|+|\nabla\phi_{\omega,j}|\leqslant C$. Thus by decreasing $\varepsilon_{8}$ again we get
$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}-\varepsilon^{2}\omega_{\mathcal{X}}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v+w)(u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j}-\varepsilon^{2}\phi_{\omega,j})\leqslant c_{X}-\delta\varepsilon+C\varepsilon^{2}<c_{X}.$
Now we fix $\varepsilon\in(0,\varepsilon_{8})$, $\sigma\in(0,\sigma_{0})$, $R>0$ and $\{B_{j,4R}\}_{j\in\mathcal{J}}$. In particular, we have a uniform bound $|\nabla\phi_{\widehat{\chi},j}|\leqslant S$ on each $B_{j,4R}$ for some $S>0$. Define
$\varphi_{j,r}:=u_{\infty,j}^{(r)}-\phi_{\widehat{\chi},j}-\varepsilon^{2}\phi_{\omega,j}$
so that
$\widehat{\chi}_{\mathcal{X}}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j,r}=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty,j}^{(r)}-\varepsilon^{2}\omega_{\mathcal{X}}>0,$
$R_{\omega_{\mathcal{X}}}\bigg(\widehat{\chi}_{\mathcal{X}}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j,r}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v+w)(\varphi_{j,r})<c_{X}.$
For $z\in B_{j,3R}$ and $r\in(0,R)$, let
$u_{j,r}(z):=\sup_{B_{j,r}(z)}u_{\infty,j},$
where $B_{j,r}(z)$ denotes a Euclidean ball of radius $r$ centered at $z$ in $B_{j,4R}$. For $r\in(0,R/2)$, we define
$\nu_{j,u_{\infty,j}}(z,r):=\frac{u_{j,\frac{3}{4}R}(z)-u_{j,r}(z)}{\log\big(\frac{3}{4}R\big)-\log r},\quad z\in B_{j,3R}.$
As $r\to 0$ the quantity $\nu_{j,u_{\infty,j}}(z,r)$ converges decreasingly to the Lelong number of $u_{\infty}$ with respect to a reference metric $\widehat{\chi}_{\mathcal{X}}$:
$\lim_{r\to 0}\nu_{j,u_{\infty,j}}(z,r)=\nu_{u_{\infty}}(z).$
Note that $\nu_{u_{\infty}}(z)=0$ for all $z\in B_{j,3R}$ since $\chi_{u_{\infty}}$ (and hence its pullback to $\mathcal{X}$) has zero Lelong numbers. By adapting the idea from Błocki–Kołodziej [BK07], Chen [Che21, Lemma 4.2] proved the following lemma, which states that the difference between $u_{j,r}(z)$ and $u_{\infty,j}^{(r)}(z)$ can be controlled in terms of $\nu_{j,u_{\infty,j}}(z,r)$.
Lemma 4.4.
For any $r\in(0,R/2)$ and $z\in B_{j,3R}$ we have
  1. (1)
    $0\leqslant u_{j,r}(z)-u_{j,\frac{r}{2}}(z)\leqslant(\log 2)\nu_{j,u_{\infty,j}}(z,r)$.
  2. (2)
    $0\leqslant u_{j,r}(z)-u_{\infty,j}^{(r)}(z)\leqslant\eta\nu_{j,u_{\infty,j}}(z,r)$, where $\eta>0$ is a universal constant depending only on $n$.
Lemma 4.5.
There exists $r_{1}\in(0,r_{0})$ such that for any $r\in(0,r_{1})$ and $z\in(B_{j,3R}\backslash B_{j,2R})\cap B_{i,R}$, we have
$\varphi_{j,r}(z)\leqslant\varphi_{i,r}(z)-\frac{\varepsilon^{2}R^{2}}{2}.$
Proof.
The function $\nu_{j,u_{\infty,j}}(z,r)$ is increasing with respect to $r$ (when fixing $z$) and upper semi-continuous with respect to $z$ (when fixing $r$). Moreover, the Lelong number $\nu_{u_{\infty}}$ is zero everywhere. Thus we can apply the Dini–Cartan lemma to know that for a constant $A>0$ determined larter, after decreasing $r_{0}$ for any $r\in(0,r_{0})$ and $z\in B_{j,3R}$ we have
$0\leqslant\nu_{j,u_{\infty,j}}(z,r)\leqslant A^{-1}\varepsilon^{2}.$
We compute
$u_{\infty,j}^{(r)}(z)$ $\leqslant u_{j,r}(z)$
$\leqslant u_{j,\frac{r}{2}}(z)+(\log 2)A^{-1}\varepsilon^{2}$
$\leqslant\sup_{B_{j,\frac{r}{2}}(z)}(\phi_{\widehat{\chi},j}+u_{\infty})+(\log 2)A^{-1}\varepsilon^{2}$
$\leqslant\sup_{B_{j,\frac{r}{2}}(z)}\phi_{\widehat{\chi},j}+\sup_{B_{j,\frac{r}{2}}(z)}u_{\infty}+A^{-1}\varepsilon^{2}.$
It follows that
$\varphi_{j,r}(z)$ $=u_{\infty,j}^{(r)}(z)-\phi_{\widehat{\chi},j}(z)-\varepsilon^{2}\phi_{\omega,j}(z)$
$\leqslant\sup_{B_{j,\frac{r}{2}}(z)}u_{\infty}+\sup_{B_{j,\frac{r}{2}}(z)}\phi_{\widehat{\chi},j}-\phi_{\widehat{\chi},j}(z)+A^{-1}\varepsilon^{2}-\varepsilon^{2}\phi_{\omega,j}(z)$
$\leqslant\sup_{B_{j,\frac{r}{2}}(z)}u_{\infty}+Sr+A^{-1}\varepsilon^{2}-3\varepsilon^{2}R^{2}.$
On the other hand,
$u_{\infty,i}^{(r)}(z)$ $\geqslant u_{i,r}(z)-\eta A^{-1}\varepsilon^{2}$
$=\sup_{B_{i,r}(z)}(\phi_{\widehat{\chi},i}+u_{\infty})-\eta A^{-1}\varepsilon^{2}$
$\geqslant\sup_{B_{i,r}(z)}u_{\infty}+\inf_{B_{i,r}(z)}\phi_{\widehat{\chi},i}-\eta A^{-1}\varepsilon^{2}.$
This shows that
$\varphi_{i,r}(z)$ $=u_{\infty,i}^{(r)}(z)-\phi_{\widehat{\chi},i}(z)-\varepsilon^{2}\phi_{\omega,i}(z)$
$\geqslant\sup_{B_{i,r}(z)}u_{\infty}+\inf_{B_{i,r}(z)}\phi_{\widehat{\chi},i}-\phi_{\widehat{\chi},i}(z)-\eta A^{-1}\varepsilon^{2}-\varepsilon^{2}\phi_{\omega,i}(z)$
$\geqslant\sup_{B_{i,r}(z)}u_{\infty}-Sr-\eta A^{-1}\varepsilon^{2}-2\varepsilon^{2}R^{2}.$
The above computation yields that
$\varphi_{i,r}(z)-\varphi_{j,r}(z)\geqslant\sup_{B_{i,r}(z)}u_{\infty}-\sup_{B_{j,\frac{r}{2}}(z)}u_{\infty}-2Sr-(\eta+1)A^{-1}\varepsilon^{2}+\varepsilon^{2}R^{2}.$
From the choice of the covering $\{B_{j,4R}\}_{j\in\mathcal{J}}$ one can easily see that $B_{j,\frac{r}{2}}(z)\subset B_{i,r}(z)$. Thus
$\varphi_{i,r}(z)-\varphi_{j,r}(z)\geqslant-2Sr-(\eta+1)A^{-1}\varepsilon^{2}+\varepsilon^{2}R^{2}.$
By choosing $A=\frac{3(\eta+1)}{R^{2}}$ we have
$\varphi_{i,r}(z)-\varphi_{j,r}(z)\geqslant-2Sr+\frac{2\varepsilon^{2}R^{2}}{3}.$
So if we set $r_{1}=\frac{\varepsilon^{2}R^{2}}{12S}$, then for any $r\in(0,r_{1})$, we conclude that
$\varphi_{j,r}(z)\leqslant\varphi_{i,r}(z)-\frac{\varepsilon^{2}R^{2}}{2}.$
The above discussion prepares the ground for constructing a subsolution on $X$. We are now ready to begin the proof.
Proof of “coercivity $\Rightarrow$ existence” in Theorem 1.2.
From Lemma 4.5, we know that for any $r\in(0,r_{1})$ the regularized maximum $\varphi$ of $\{(B_{j,3R},\varphi_{j,r})\}_{j\in\mathcal{J}}$ (for a sufficiently small vector $\eta$) is smooth on $\mathcal{X}_{1}$. Define $\Omega:=\widehat{\chi}_{\mathcal{X}}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi$. Then by Proposition 2.13, $\Omega$ is a Kähler form on $\mathcal{X}_{1}$ which satisfies
$R_{\omega_{\mathcal{X}}}(\Omega)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v+w)(\varphi)<c_{X}.$
Let
$\widetilde{\varphi}:=\int_{\tau\in T^{\prime}}\tau^{\ast}\varphi dT^{\prime},\quad\widetilde{\Omega}=\widehat{\chi}_{\mathcal{X}}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\widetilde{\varphi}$.
[…]
$R_{\omega_{\mathcal{X}}}(\widetilde{\Omega})-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\widetilde{\varphi})<c_{X}$
on $\mathcal{X}_{1}$. Finally, we observe that the restriction of $\widetilde{\Omega}$ to $X\simeq X\times\{1\}\subset\mathcal{X}_{1}$:
$\chi:=\widetilde{\Omega}|_{X}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\widetilde{\varphi}|_{X}$
satisfies
$P_{\omega}(\chi)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\widetilde{\varphi}|_{X})<c_{X},$
where we used Proposition 2.2 and the fact that $\mathrm{Re}(v)$ is tangent to $X$. Hence, by Proposition 2.15, if we set
$\phi:=\int_{\tau\in T}\tau^{\ast}\widetilde{\varphi}|_{X}dT,\quad\widetilde{\chi}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi,$
then the Kähler form $\widetilde{\chi}\in\beta$ satisfies
$P_{\omega}(\widetilde{\chi})-\theta_{v}^{X}(\widetilde{\chi})<c_{X}.$
The desired result follows from Theorem 1.1. ∎
4.2. Existence implies coercivity
We will prove the converse direction of Theorem 1.2 and Theorem 1.3. As in Proposition 4.1, the terms involving the Hamiltonian cancel out when taking the difference of the functionals, so that, in effect, there is no substantial difference from the usual $J$-equation case [CS17, Proposition 22]. However, we will give it here for the sake of completeness.
Proof of “existence $\Rightarrow$ coerxivity” in Theorem 1.2.
Assume that there exists a solution $\widehat{\chi}\in\beta$ to the modified $J$-equation
[…]
Take a sufficiently small $\delta>0$ so that $\omega^{\prime}:=\omega-\delta\widehat{\chi}>0$. Then
[…]
i.e. the Kähler form $\widehat{\chi}$ solves the modified $J$-equation with respect to $\omega^{\prime}$. In particular, we have $J_{v}^{\omega^{\prime}}(\phi)\geqslant-C$ for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ [LS16, Corollary 3.2]. For all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, we compute
$J_{v}^{\omega}(\phi)=J_{v}^{\omega^{\prime}}(\phi)-n\delta\int_{0}^{1}\int_{X}\phi(\chi_{t\phi}^{n}-\widehat{\chi}\wedge\chi_{t\phi}^{n-1})dt.$
Integrating by parts we get
$-\int_{X}\phi(\chi_{t\phi}^{n}-\widehat{\chi}\wedge\chi_{t\phi}^{n-1})$ $=-\int_{X}\phi\bigg(t\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi\bigg)\wedge\chi_{t\phi}^{n-1}$
$=\frac{\sqrt{-1}}{2\pi}t\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge\chi_{t\phi}^{n-1}$
$=\frac{\sqrt{-1}}{2\pi}\sum_{i=1}^{n-1}p_{i}(t)\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge\chi_{\phi}^{i}\wedge\widehat{\chi}^{n-1-i},$
where $p_{i}(t):=\binom{n-1}{i}t^{i+1}(1-t)^{n-1-i}$. We take a constant $\kappa>0$ (depending only on $n$) sufficiently small so that
$\int_{0}^{1}p_{i}(t)dt\geqslant\kappa,\quad i=1,\ldots n-1.$
Then we have
$-n\delta\int_{0}^{1}\int_{X}\phi(\chi_{\phi_{t}}^{n}-\widehat{\chi}\wedge\chi_{\phi_{t}}^{n-1})dt$ $=\frac{\sqrt{-1}}{2\pi}n\delta\sum_{i=1}^{n-1}\int_{0}^{1}p_{i}(t)dt\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge\chi_{\phi}^{i}\wedge\widehat{\chi}^{n-1-i}$
$\geqslant\frac{\sqrt{-1}}{2\pi}n\delta\kappa\int_{X}\partial\phi\wedge\bar{\partial}\phi\wedge\sum_{i=1}^{n-1}\chi_{\phi}^{i}\wedge\widehat{\chi}^{n-1-i}$
$=n\delta\kappa I(\phi).$
It then follows that
$J_{v}^{\omega}(\phi)\geqslant n\delta\kappa I(\phi)-C$
for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$. ∎
As a corollary of this we can show that the solvability of the $J$-equation does not depend on a choice of $\omega\in\alpha$.
Proof of Theorem 1.3.
Thanks to Theorem 1.2, it suffices to show that the coercivity of $J_{v}^{\omega}$ does not depend on the choice of $\omega$. If $\omega,\omega^{\prime}\in\alpha$ are $T$-invariant Kähler forms there exists a $T$-invariant smooth function $\varphi_{0}$ such that $\omega=\omega^{\prime}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{0}$. Then we compute
$J_{v}^{\omega}(\phi)-J_{v}^{\omega^{\prime}}(\phi)$ $=n\int_{0}^{1}\int_{X}\phi\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{0}\wedge\chi_{t\phi}^{n-1}dt$
$=n\int_{0}^{1}\int_{X}\varphi_{0}\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi\wedge\chi_{t\phi}^{n-1}dt$
$=\int_{0}^{1}\int_{X}\varphi_{0}\frac{d\chi_{t\phi}^{n}}{dt}dt$
$=\int_{X}\varphi_{0}(\chi_{\phi}^{n}-\widehat{\chi}^{n}),$
and hence we have
$\big|J_{v}^{\omega}(\phi)-J_{v}^{\omega^{\prime}}(\phi)\big|\leqslant\int_{X}|\varphi_{0}|\chi_{\phi}^{n}+C\leqslant C$
as desired. ∎
5. A Nakai–Moishezon type criterion for smooth projective toric varieties
5.1. Local $C^{2}$-estimate for the generalized equation
Let $X$ be a compact complex manifold, $T\subset\operatorname{Aut}_{\rm red}(X)$ a real torus, and let $\omega\in\alpha$ and $\widehat{\chi}\in\beta$ be $T$-invariant Kähler forms. Let $b\geqslant 0$ and $c>0$ be constants related by (1.7). In this subsection we establish the local $C^{2}$-estimate of the generalized equation for $\chi_{\phi}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi$ ($\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$):
$\operatorname{Tr}_{\chi_{\phi}}\omega+b\frac{\omega^{n}}{\chi_{\phi}^{n}}=c+\theta_{v}^{X}(\chi_{\phi}).$ (5.1)
Lemma 5.1.
Let $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ be a solution to (5.1). Assume that there exists a Kähler form $\widetilde{\chi}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\psi$ defined on the closure $\overline{\Omega}$ of an open subset $\Omega\subset X$ satisfying
$P_{\omega}(\widetilde{\chi})-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\psi)<c-2\delta$
for some $\delta>0$ (where $\Omega$, $\psi$ and $\widetilde{\chi}$ do not have to be $T$-invariant). Then
$\operatorname{Tr}_{\omega}\chi_{\phi}\leqslant Ce^{N(\phi-\inf_{\overline{\Omega}}\phi)}$
on $\overline{\Omega}$, where the constants $C>0$, $N>0$ depend on $\widetilde{\chi}$, $\psi$, $\delta$, $C_{\theta}$, an upper bound of $c$ and the maximum of $\operatorname{Tr}_{\omega}\chi_{\phi}$ on $\partial\Omega$.
As a concrete situation in which the above Lemma is applied, we consider $X$ to be toric and $\Omega$ an open subset of $X$ containing the union of all toric divisors. This Lemma is required for the inductive proof of Theorem 5.2 on toric subvarieties.
Proof of Lemma 5.1.
The proof is essentially the same as Lemma 3.4. Set $\varphi:=\phi-\psi$ so that $\chi_{\phi}=\widetilde{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi$. Define the operator $\mathcal{L}_{b}$ by
$\mathcal{L}_{b}:=\partial_{ij}F_{b}(A)g^{i\bar{k}}\nabla_{j}\nabla_{\bar{k}}-\mathrm{Re}(v),$
where $A=g^{i\bar{k}}h_{j\bar{k}}$ and $\nabla$ denotes the Chern connection with respect to $\omega$. We take normal coordinates of $\omega$ on which $\chi_{\phi}$ is diagonal. From the equation (5.1) we know that $\chi_{\phi}\geqslant C_{1}^{-1}\omega$ for some constant $C_{1}>0$ depending on $C_{\theta}$ and an upper bound of $c$. Then we have
$\mathcal{L}_{b}\log\operatorname{Tr}_{\omega}\chi_{\phi}\leqslant C_{2}$
(indeed, we may substitute $\frac{d}{dt}\phi_{t}=0$ to the computation in Lemma 3.4 since $\phi$ is a stationary point of the modified $J$-flow). Similarly by using (5.1) and $\varphi=\phi-\psi$ a direct computation shows that
$\mathcal{L}_{b}\varphi=-F_{b}(A)+c+\theta_{v}^{X}(\widehat{\chi})+\mathrm{Re}(v)(\psi)+\partial_{ii}F_{b}(A)(h_{i\bar{i}}-\widetilde{h}_{i\bar{i}}),$
where $\widetilde{h}$ denotes the Riemannian metric corresponding to $\widetilde{\chi}$. Define the function $G\colon\overline{\Omega}\to\mathbb{R}$ by $G:=\log\operatorname{Tr}_{\omega}\chi_{\phi}-N\varphi$. Then the above computation shows that
$\mathcal{L}_{b}G\leqslant C_{2}+NF_{b}(A)-Nc-N\theta_{v}^{X}(\widehat{\chi})-N\mathrm{Re}(v)(\psi)+N\partial_{ii}F_{b}(A)(\widetilde{h}_{i\bar{i}}-h_{i\bar{i}}).$
We take a large constant $N>0$ so that $C_{2}\leqslant N\delta$. If the function $G$ takes maximum at some point $x_{0}\in\Omega$ then we have
$\delta\leqslant F_{b}(A)-c-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\psi)+\partial_{ii}F_{b}(A)(\widetilde{h}_{i\bar{i}}-h_{i\bar{i}}),$
and hence
$P_{\omega}(\widetilde{\chi})+\delta\leqslant F_{b}(A)+\partial_{ii}F_{b}(A)(\widetilde{h}_{i\bar{i}}-h_{i\bar{i}})$
at $x_{0}$. Thus we can apply the same argument in the second half of the proof for Lemma 5.1 just by replacing $\widehat{\chi}$ with $\widetilde{\chi}$, and get
$\operatorname{Tr}_{\omega}\chi_{\phi}\leqslant Ce^{N(\varphi-\inf_{\overline{\Omega}}\varphi)}.$
Finally, by using a bound of $\psi$, we obtain the desired estimate. ∎
5.2. Continuity method and induction for toric subvarieties
In what follows let $X$ be an $n$-dimensional smooth projective toric variety associated to a fan $\Sigma$, and let $T\subset T^{\mathbb{C}}\simeq(\mathbb{C}^{\ast})^{n}$ denote the real torus and its complexification associated with $X$. We denote $N:=\operatorname{Hom}(\mathbb{C}^{\ast},T^{\mathbb{C}})$ the lattice of one-parameter subgroups of $T^{\mathbb{C}}$. The fan $\Sigma$ is then a collection of strictly convex rational polyhedral cones in $N_{\mathbb{R}}:=N\otimes_{\mathbb{Z}}\mathbb{R}$.
Let $v$ be a holomorphic vector field with $\mathrm{Im}(v)\in\mathfrak{t}$, and assume that Kähler forms $\omega\in\alpha$, $\widehat{\chi}\in\beta$ are $T$-invariant. For any $p$-dimensional toric subvarieties [^1] In this paper the term “$p$-dimensional subvariety” means that it is reduced, but not necessarily irreducible or purely $p$-dimensional. $Y\subset X$ we define the set $\Gamma_{\omega,\widehat{\chi}}(Y)$ consisting of germs of $\widehat{\chi}-$Kähler potential $\phi$ satisfying
$Q_{X,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\phi)<c_{X}$
at $Y\subset X$. When $Y$ is irreducible, it is given as the $T^{\mathbb{C}}$-invariant subvariety associated with a cone $\tau\in\Sigma$
$Y=\overline{O(\tau)}\subset X,$
which is the closure of the $T^{\mathbb{C}}$-orbit corresponding to $\tau$. Let $N_{\tau}$ be the sublattice of $N$ spanned by the points in $N\cap\tau$ and define $N(\tau)$ by the exact sequence
$0\to N_{\tau}\to N\to N(\tau)\to 0.$
Then the fan $\Sigma(\tau)$ corresponding to $Y$ is given by
$\Sigma(\tau)=\{\overline{\sigma}\subset N(\tau)_{\mathbb{R}}|\tau\preceq\sigma\in\Sigma\},$
where $\overline{\sigma}$ denotes the image cone of $\sigma$ in $N(\tau)_{\mathbb{R}}:=N(\tau)\otimes_{\mathbb{Z}}\mathbb{R}$ under the quotient map $N_{\mathbb{R}}\to N(\tau)_{\mathbb{R}}$. In particular, the properties of a cone, rationality, smoothness and strict convexity are inherited by the toric subvariety $Y$. Also the real torus associated with $Y$ is then given by $T_{Y}:=N(\tau)\otimes_{\mathbb{Z}}\mathbb{S}^{1}$. Thus tensoring with $\mathbb{S}^{1}$ to the projection $N\to N(\tau)$, we obtain a surjective homomorphism $T\to T_{Y}$. The $T$-action on $Y$ is just obtained by composing this morphism with the $T_{Y}$-action on $Y$ (for the notation and a more detailed explanation, see [CLS11, Section 3.2]). Define
$\mathcal{H}(Y,\widehat{\chi})^{T_{Y}}:=\big\{\phi\in C^{\infty}(Y;\mathbb{R})^{T_{Y}}\big|\chi_{\phi}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi>0\big\}.$
[…]
$\theta_{v}^{Y}(\widehat{\chi})=\theta_{v}^{X}(\widehat{\chi})-I_{Y},$
where
$I_{Y}:=\frac{1}{\beta^{p}}\int_{Y}\theta_{v}^{X}(\widehat{\chi})\widehat{\chi}^{p},$
and we often omit the notation for restriction to $Y$. Since $\theta_{v}^{Y}(\chi_{\phi})=\theta_{v}^{Y}(\widehat{\chi})+v(\phi)$ we have
$\theta_{v}^{Y}(\chi_{\phi})=\theta_{v}^{X}(\widehat{\chi})+v(\phi)-I_{Y}$ (5.2)
for all $\phi\in\mathcal{H}(Y,\widehat{\chi})^{T_{Y}}$.
Theorem 5.2.
Let $X$, $T$, $v$, $\omega$, $\widehat{\chi}$, $\alpha$, $\beta$ as above. Assume $m_{X}>0$ and there exists a $T$-invariant Kähler form $\chi\in\beta$ such that
$\int_{Y}\big((c_{X}+\theta_{v}^{X}(\chi))\chi^{p}-p\omega\wedge\chi^{p-1}\big)>0$
for all $p$-dimensional irreducible toric subvarieties $Y\subset X$ ($p=1,\ldots,n-1$). Then for any $p$-dimensional toric subvarieties $Y\subset X$ ($p=0,1,\ldots,n-1$) we have
$\Gamma_{\omega,\widehat{\chi}}(Y)\neq\emptyset.$
Proof.
We will prove by induction of $p=\dim Y$.
Step 0. First, we note that if $\dim Y=0$, i.e. $Y$ is a finite set of points, then the result easily follows from $m_{X}=c_{X}+\min_{X}\theta_{v}^{X}(\widehat{\chi})>0$, and the facts that $\mathrm{Re}(v)$ vanishes at $Y$ and every Kähler class in a sufficiently small neighborhood of a point is trivial.
Step 1. Assume that $p\geqslant 1$ and the statement is true for all toric subvarieties $Z\subset X$ of $\dim Z\leqslant p-1$. For any $p$-dimensional irreducible toric subvariety $Y\subset X$ let us consider the following continuity path $\chi_{\phi_{t}}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{t}\in\beta|_{Y}$ ($t\in[0,\infty)$) with $\phi_{t}\in\mathcal{H}(Y,\widehat{\chi})^{T_{Y}}$:
$\operatorname{Tr}_{\chi_{\phi_{t}}}\omega+b_{t}\frac{\omega^{p}}{\chi_{\phi_{t}}^{p}}=c_{t}+\theta_{v}^{Y}(\chi_{\phi_{t}}),$ (5.3)
where
$c_{t}:=c_{X}(t+1)+I_{Y}\geqslant c_{X}+I_{Y}\geqslant c_{X}+\min_{X}\theta_{v}^{X}(\widehat{\chi})=m_{X}>0,$
$b_{0}:=\frac{(c_{X}+I_{Y})\beta^{p}-p\alpha\cdot\beta^{p-1}}{\alpha^{p}}>0,$
$b_{t}:=\frac{c_{t}\beta^{p}-p\alpha\cdot\beta^{p-1}}{\alpha^{p}}=b_{0}+c_{X}\frac{\beta^{p}}{\alpha^{p}}t>0$
from the assumption. We note that
[…]
$\geqslant c_{X}+I_{Y}+\min_{Y}(\theta_{v}^{X}(\widehat{\chi})-I_{Y})$
[…]
$=m_{X}>0,$
which will be required to obtain the interior $C^{2}$-estimates (see Section 5.3). Set
[…]
By Theorem 1.1 the solvability of (5.3) is equivalent to the existence of $\underline{\phi}\in\mathcal{H}(Y,\widehat{\chi})^{T_{Y}}$ such that
$P_{\omega}(\chi_{\underline{\phi}})-\theta_{v}^{Y}(\chi_{\underline{\phi}})<c_{t}.$ (5.4)
From the uniform bound of $\theta_{v}^{Y}(\chi_{\underline{\phi}})$, we know that any $\underline{\phi}\in\mathcal{H}(Y,\widehat{\chi})^{T_{Y}}$ satisfies (5.4) for sufficiently large $t$. It then follows that $\mathcal{T}$ is open, and there exists a large constant $T>0$ such that $T,\infty)\subset\mathcal{T}$.
Now we will show that $\mathcal{T}$ is closed. Let $t_{0}:=\inf\mathcal{T}$ and $D=\cup_{i}D_{i}$ be the union of all toric divisors of $Y$. From the inductive hypothesis, there exists an open neighborhood $U_{D}$ of $D$ in $X$, and a $\widehat{\chi}-$Kähler potential $\phi_{D}$ on $\overline{U_{D}}$ such that
$Q_{X,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{D}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\phi_{D})<c_{X}-\delta$
for some $\delta>0$. By Proposition (2.2) (1) and (5.2), we have
$Q_{Y,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{D}\bigg)-\theta_{v}^{Y}(\widehat{\chi})-\mathrm{Re}(v)(\phi_{D})<c_{X}+I_{Y}-\delta\leqslant c_{t_{0}}-\delta$
on the closure $\overline{U_{Y}}$ of $U_{Y}:=U_{D}\cap Y$ in $Y$, since $\mathrm{Re}(v)$ is tangent to $Y$. For $t>t_{0}$ close to $t_{0}$, we invoke Lemma 5.1 to reduce the $C^{2}$-estimates of $\phi_{t}$ on $\overline{U_{Y}}$ to estimates on the boundary $\partial U_{Y}$. However, by Proposition 5.4 we obtain bounds of the form
$\operatorname{Tr}_{\omega}\chi_{\phi_{t}}\leqslant Ce^{N(\phi_{t}-\inf_{Y\backslash U_{Y}}\phi_{t})}$
on $Y\backslash U_{Y}$. Eventually, we obtain the same inequality globally on $Y$. Thus we can use the argument [SW08, Wei04, Wei06], the Evans–Klyrov theorem and the Schauder estimates to obtain higher order estimates. This shows that $\mathcal{T}$ is closed, and hence $\mathcal{T}=[0,\infty)$.
Step 2. By Step 1 we have obtained $\phi\in\mathcal{H}(Y,\widehat{\chi})^{T_{Y}}$ satisfying
$Q_{Y,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\phi)\leqslant c_{X}-2\delta$
as a solution to (5.3) for $t=0$ on $Y$, where the existence of a positive constant $\delta>0$ follows from $b_{0}>0$. For a projective toric variety $X$, the rational, complete fan $\Sigma$ and lattice $N$ allow a canonical $T_{Y}$-equivariant linearization on the line bundle $\mathcal{O}_{Y}(D_{i})$ [CLS11, Section 6] [^2] In general, for a smooth projective variety $X$, a real torus $T\subset\operatorname{Aut}_{\rm red}(X)$ and a line bundle $L$, the bundle $L$ admits a $T$-equivariant linerization (cf. [Kob95, Chapter III, Proposition 9.5]). However, when $X$ is not projective, the existence of a $T$-equivariant linearization on $\mathcal{O}_{X}(D)$ is obstructed even if the divisor $D\subset X$ is $T$-invariant [KS24, Theorem 2].. We take a defining section $s_{i}$ of $D_{i}$ and $T_{Y}$-invariant fiber metric $h_{i}$ on $\mathcal{O}_{Y}(D_{i})$. Then the curvature $\xi_{i}$ of $h_{i}$ has an expression
$\xi_{i}=-\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\log|s_{i}|_{h_{i}}^{2}$ (5.5)
on $Y\backslash D_{i}$. In particular we have a uniform bound
$C\omega\leqslant\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\log|s_{i}|_{h_{i}}^{2}\leqslant C\omega$
on $Y\backslash D_{i}$. For any $t\in T_{Y}$, both $s_{i}$ and $t\cdot s_{i}$ are defining sections of $D_{i}$, which implies that $t\cdot s_{i}=\chi(t)s_{i}$ for some character $\chi\colon T_{Y}\to\mathbb{C}^{\ast}$. Since $T_{Y}$ is compact, the image of $\chi$ must be a compact subgroup of $\mathbb{C}^{\ast}$, which shows that $|\chi(t)|=1$ for all $t\in T_{Y}$. In particular, we know that $|s_{i}|_{h}^{2}$ is $T_{Y}$-invariant and $\mathrm{Im}(v)\log|s_{i}|_{h}^{2}=0$ on $Y\backslash D_{i}$. Applying $i_{v}$ to (5.5) we get
$i_{v}\xi_{i}=-\frac{\sqrt{-1}}{2\pi}\bar{\partial}\big(\mathrm{Re}(v)(\log|s_{i}|_{h_{i}}^{2})\big),$ (5.6)
and hence
$\big|\mathrm{Re}(v)(\log|s_{i}|_{h_{i}}^{2})\big|\leqslant C$
on $Y\backslash D_{i}$. Thus for sufficiently small $\kappa>0$ the function
$\psi:=\phi+\kappa\sum_{i}\log|s_{i}|_{h_{i}}^{2}$
satisfies
$Q_{Y,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\psi\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\psi)\leqslant c_{X}-\delta$
[…]
$Q_{X,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{Z}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\phi_{Z})<c_{X}$
on $U$. On the other hand, applying the above construction to each $Y_{k}$, we obtain $\psi_{k}$ such that
$Q_{Y_{k},\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\psi_{k}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\psi_{k})\leqslant c_{X}-\delta$
on $Y_{k}^{\prime}$ and $\psi_{k}\to$-$\infty$ along the toric divisors of $Y_{k}$. For simplicity, we assume that $\ell=1$ (the general case follows by applying the same argument to each $Y_{k}$). We can choose open neighborhoods $W\Subset V\Subset U$ of $Z$ in $X$ such that
  1. (1)
[…]
  2. (2)
    $\psi_{1}<\phi_{Z}-2$ in $Y_{1}\cap W$
[…]
$\widetilde{\psi}_{1}:=\psi_{1}+Bd_{Y_{1}}^{2},$
where $B>0$ is a constant and $d_{Y_{1}}$ is the distance function from $Y_{1}$. By taking $B$ sufficiently large, we have
$Q_{X,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\widetilde{\psi}_{1}\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\widetilde{\psi}_{1})<c_{X}$
[…]
$\widetilde{\psi}_{1}>\phi_{Z}+1$
[…]
$\widetilde{\psi}_{1}<\phi_{Z}-1$
in $\widetilde{U}\cap W$. Let $\varphi$ be the regularized maximum of $(\widetilde{U},\widetilde{\psi}_{1})$ and $(V,\phi_{Z})$. From the above observations, we conclude that
$Q_{X,\omega}\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi\bigg)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\varphi)<c_{X}$
on a small neighborhood of $Y$ in $X$. This completes the proof. ∎
Proof of Theorem 1.6.
In Step 1 of Theorem 5.2, let $Y=X$ and $D$ the union of all toric divisors of $X$. Then by Theorem 5.2, we know that $\Gamma_{\omega,\widehat{\chi}}(D)\neq\emptyset$. Hence by proceeding with the same argument, we see that the continuity path
$\operatorname{Tr}_{\chi_{\phi_{t}}}\omega+b_{t}\frac{\omega^{n}}{\chi_{\phi_{t}}^{n}}=c_{t}+\theta_{v}^{X}(\chi_{\phi_{t}}),\quad b_{t}=c_{X}\frac{\beta^{n}}{\alpha^{n}}t,\quad c_{t}=c_{X}(t+1)$
for $\chi_{\phi_{t}}=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi_{t}\in\beta$ ($\phi_{t}\in\mathcal{H}(X,\widehat{\chi})^{T}$) has a solution for all $t\in[0,\infty)$. This completes the proof. ∎
5.3. Interior $C^{2}$-estimates
The goal of this subsection is to prove the interior $C^{2}$-estimates required in Step 1 of Theorem 5.2. Let $X$ be an $n$-dimensional smooth projective toric variety. We identity the open dense $T$-orbit in $X$ with $\mathbb{R}^{n}\times(\mathbb{S}^{1})^{n}$ and express $\omega$, $\chi$ as
$\omega=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}f,\quad\chi=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}g$
for some smooth strictly convex function $f,g\colon\mathbb{R}^{n}\to\mathbb{R}$. Assume that $f$, $g$ satisfy the equation
$g^{ij}f_{ij}+b\frac{\det(D^{2}f)}{\det(D^{2}g)}=c+\ell(\nabla g)-\frac{1}{V}\int_{P}\ell dy,$ (5.7)
where $b\geqslant 0$ and $c>0$ are constants, $V:=\operatorname{Vol}(P)$, $\ell\colon\mathbb{R}^{n}\to\mathbb{R}$ is a linear function, and $P=\nabla g(\mathbb{R}^{n})$ is the moment polytope determined uniquely from the Kähler class $\beta$ up to parallel translations. Also we assume that
$c+\inf_{P}\ell-\frac{1}{V}\int_{P}\ell dy\geqslant\delta$ (5.8)
for some constant $\delta>0$. Adding an affine linear function to $g$ amounts to a translation of the moment polytope $P$. Since the equation (5.7) and the condition (5.8) remain unchanged under this operation, we may assume that $g(0)=0$ and $\nabla g(0)=0$. In particular, since $0\in P$ we obtain:
Lemma 5.3 ([CS17], Lemma 26).
For any compact subset $K\subset\mathbb{R}^{n}$ there exists $C>0$ such that $\sup_{K}|g|<C$.
Let $h\colon P\to\mathbb{R}$ be the Legendre transform of $g$. Then we observe that
$\nabla h(y)=(\nabla g)^{-1}(y),\quad D^{2}h(y)=\big(D^{2}g(\nabla h(y))\big)^{-1}.$ (5.9)
By using (5.9) we can translate the equation (5.7) to
$f_{ij}(\nabla h(y))h_{ij}+b\det(D^{2}f(\nabla h(y)))\det(D^{2}h(y))=c+\ell(y)-\frac{1}{V}\int_{P}\ell dy.$ (5.10)
Our strategy is essentially the same as [CS17, Section 5.1], namely, we establish higher order estimates for the Legendre transform $h$. If there is a sequence of equations (5.7) that violates the $C^{2}$ bound, we take the limit $k\to\infty$ and apply Bian–Guan’s constant rank theorem [BG09, Theorem 1.1] to obtain a contradiction.
Proposition 5.4.
Let $B\subset\mathbb{R}^{n}$ be the unit ball, and $f$, $g$ satisfy (5.7) with normalization $\inf_{B}g=g(0)=0$. Then there exists $C>0$ depending only on $\sup_{B}|g|$, bounds on $b$, $c$, the coefficients of $\ell$, $C^{3,\gamma}$ bounds on $f$ and a positive lower bound on the Hessian of $f$ such that
$\sup_{\frac{1}{2}B}|g_{ij}|<C.$
Proof.
Suppose that there are sequences $f_{k}$, $g_{k}$, $b_{k}$, $c_{k}$, $\ell_{k}$ satisfying the hypothesis, including $|g_{k}|<N$ for some constant $N>0$, but $|D^{2}g_{k}(x_{k})|>k$ for some $x_{k}\in\frac{1}{2}B$. The moment polytopes $P_{k}=\nabla g_{k}(\mathbb{R}^{n})$ of $g_{k}$ satisfy $0\in P_{k}$ and are parallel to each other. So by passing to a subsequence we may assume that $P_{k}\to P_{\infty}\subset\mathbb{R}^{n}$ via parallel translations, and hence $\int_{P_{k}}\ell_{k}dy\to\int_{P_{\infty}}\ell_{\infty}dy$ as $k\to\infty$. By shrinking the ball a bit we can assume that $g_{k}\to g$ uniformly for some strictly convex $g\colon B\to\mathbb{R}$. Also, by (5.7) we have
$g_{k}{}^{ij}f_{k,ij}\leqslant c_{k}+\max_{P_{k}}\ell_{k}-\frac{1}{V}\int_{P_{k}}\ell_{k}dy.$
Since the right hand side has a uniform upper bound, combining with a uniform lower bound for $D^{2}f_{k}$ yields
$g_{ij}>\tau\delta_{ij}$ (5.11)
for some $\tau>0$ independent of $k$. Let $h_{k}$ be the Legendre transform of $g_{k}$ and set $U_{k}:=(\nabla g_{k})(B)\subset P_{k}$. Then we have $h_{k}(0)=0$, $\nabla h_{k}(0)=0$ from the normalization. Moreover, the formula (5.9) together with (5.11) yields an upper bound of $D^{2}h_{k}$, and hence $h_{k}$ is uniformly bounded in $C^{2}$. Let $\nabla g$ denote the subdifferential of $g$, i.e.
[…]
and
$(\nabla g)(E):=\bigcup_{x\in E}(\nabla g)(x)$
for any subset $E\subset B$. Then the estimate for subdifferentials [CS17, Lemma 28] implies that for sufficiently large $k$ we have $(\nabla g)(0.9B)\subset U_{k}$, and so $(\nabla g)(0.8B)$ remains a fixed positive distance from $\partial U_{k}$ for large $k$. In addition $h_{k}$ solves the equation
$f_{k,ij}(\nabla h_{k}(y))h_{k,ij}+b_{k}\det(D^{2}f_{k}(\nabla h_{k}(y)))\det(D^{2}h_{k}(y))=c_{k}+\ell_{k}(y)-\frac{1}{V}\int_{P_{k}}\ell_{k}dy.$
We use Proposition 5.5 together with the Schauder estimate, to obtain uniform $C^{3,\gamma}$ bounds for $h_{k}$ on $\nabla g(0.8B)$. So combining with bounds for $f_{k}$, $b_{k}$, $c_{k}$, $\ell_{k}$, we can extract a convergent subsequence $h_{k}\to h_{\infty}$ on $\nabla g(0.8B)$ in $C^{3,\gamma/2}$ satisfying an equation of the form
$f_{\infty,ij}(\nabla h_{\infty}(y))h_{\infty,ij}+b_{\infty}\det(D^{2}f_{\infty}(\nabla h_{\infty}(y)))\det(D^{2}h_{\infty}(y))=c_{\infty}+\ell_{\infty}(y)-\frac{1}{V}\int_{P_{\infty}}\ell_{\infty}dy.$
As in [CS17, Proposition 27] we consider the following two cases separately:
  • If we have a positive lower bound on $D^{2}h_{\infty}$ this yields a uniform lower bound on $D^{2}h_{k}$ on $\nabla g(0.8B)$, and so a uniform upper bound for $D^{2}g_{k}$ at all points for which $\nabla g_{k}(x)\in\nabla g(0.8B)$. However, we have $\nabla g_{k}(0.7B)\subset\nabla g(0.8B)$ for sufficiently large $k$ by [CS17, Lemma 28], and hence an upper bound of $D^{2}g_{k}$ on $0.7B$. This is a contradiction.
  • If the Hessian $D^{2}h_{\infty}$ degenerates as some point then we apply the Bian–Guan’s constant rank theorem [BG09, Theorem 1.1] to know that $D^{2}h_{\infty}$ degenerates everywhere. Indeed, for a fixed value of $\nabla h_{\infty}$, the equation is of the form
    $F(A,y):=\operatorname{Tr}(BA)+b\det(A)-L(y)=0$
    for a positive definite Hermitian matrix $B$, constant $b\geqslant 0$ and affine linear function $L$. The assumptions of [BG09, Theorem 1.1] are satisfied, since $F(A^{-1},y)$ is convex in $(A,y)$ according to Proposition 2.9. Thus we have
    $\int_{\nabla g(0.8B)}\det(D^{2}h_{\infty})=0.$
    On the other hand,
    $\int_{\nabla g_{k}(0.7B)}\det(D^{2}h_{k})=\operatorname{Vol}(0.7B).$
    This contradicts the fact that $\nabla g_{k}(0.7B)\subset\nabla g(0.8B)$ for sufficiently large $k$ and $h_{k}\to h_{\infty}$ in $C^{3,\gamma/2}$.
Hence, the local estimate for $D^{2}g$ can be reduced to the local $C^{2,\gamma}$-estimate for $h$, which we will focus in the following. A main difference from the usual $J$-equation case [CS17, Proposition 29] is that we have to deal with an affine linear function in the right hand side of (5.10). However we will observe that the affine linear function converges to a positive constant through the blowup argument. So the Liouville rigidity result [CS17, Lemma 31] can also be applied to our case.
Proposition 5.5.
Suppose that $h$ is a smooth convex function on the unit ball $B\subset\mathbb{R}^{n}$ satisfying the equation
$a_{ij}(\nabla h)h_{ij}+b(\nabla h)\det(D^{2}h)=L,$ (5.12)
where $a_{ij},b\in C^{1,\gamma}$, $b\geqslant 0$, $\lambda<a_{ij}<\Lambda$ for some constants $\lambda,\Lambda>0$, $L$ is an affine linear function with $\inf_{B}L\geqslant\delta$ for some constant $\delta>0$. Then there exists a constant $C>0$ depending on $\lambda$, $\Lambda$, $\delta$, bound of the coefficients of $L$, $C^{1,\gamma}$ bounds of $a_{ij}$ and $b$, and the $C^{2}$ bounds of $h$, such that
$\|h\|_{C^{2,\gamma}(\frac{1}{2}B)}<C.$
Proof.
Let
$N_{h}:=\sup_{y\in B}d_{y}|D^{3}h(y)|,$
where $d_{y}:=d(y,\partial B)$ denotes the distance from the boundary $\partial B$. Assume that $N_{h}>1$ and the supremum is achieved at a point $y_{0}\in B$. Define the function
$\widetilde{h}(y)=d_{y_{0}}^{-2}N_{h}^{2}h(y_{0}+d_{y_{0}}N_{h}^{-1}y)-\widetilde{L}(y),$
where the affine function $\widetilde{L}$ is chosen so that
$\widetilde{h}(0)=0,\quad\nabla\widetilde{h}(0)=0.$ (5.13)
Note that the rescaled function $\widetilde{h}$ is defined on the ball $B_{N_{h}}(0)$. Also a direct computation shows that
$D^{2}\widetilde{h}(y)=D^{2}h(y_{0}+d_{y_{0}}N_{h}^{-1}y),\quad D^{3}\widetilde{h}(y)=d_{y_{0}}N_{h}^{-1}D^{3}h(y_{0}+d_{y_{0}}N_{h}^{-1}y).$
In particular, we have $|D^{3}\widetilde{h}(y)|\leqslant 2$ on $B_{2^{-1}N_{h}}(0)$. Moreover, since $|D^{2}\widetilde{h}|=|D^{2}h|<C$ the normalization (5.13) implies that
$\|\widetilde{h}\|_{C^{3}(B_{2^{-1}N_{h}}(0))}<C$
for a uniform constant $C>0$. One can observe that the rescaled function $\widetilde{h}$ satisfies an equation of the form
$\widetilde{a}_{ij}(\nabla\widetilde{h}(y))\widetilde{h}_{ij}(y)+\widetilde{b}(\nabla\widetilde{h}(y))\det(D^{2}\widetilde{h}(y))=L(y_{0}+d_{y_{0}}N_{h}^{-1}y),$ (5.14)
where the coefficients $\widetilde{a}_{ij}$, $\widetilde{b}$ has the same $C^{0}$ bounds as $a_{ij}$, $b$, whereas
$\sup|\nabla\widetilde{a}_{ij}|\leqslant d_{y_{0}}N_{h}^{-1}\sup|\nabla a_{ij}|,\quad\sup|\nabla\widetilde{b}|\leqslant d_{y_{0}}N_{h}^{-1}\sup|\nabla b|.$ (5.15)
The differentiating the equation (5.14) and using the standard Schauder theory we know that $\nabla\widetilde{h}$ is bounded in $C^{2,\gamma}$.
Now we argue by contradiction. So we suppose that there exists a sequence of convex functions $h_{k}$ on $B$ satisfying (5.12) with coefficients $a_{k,ij}$, $b_{k}$, and $L_{k}$, and associated constant $N_{h_{k}}>4k$. For each $k$ we take a point $y_{0,k}\in B$ that attains the supremum of $N_{h_{k}}$. Then the rescaled function $\widetilde{h}_{k}$ defined on $B_{4k}(0)$ satisfies an equation of the form
$\widetilde{a}_{k,ij}(\nabla\widetilde{h}_{k}(y))\widetilde{h}_{k,ij}(y)+\widetilde{b_{k}}(\nabla\widetilde{h}_{k}(y))\det(D^{2}\widetilde{h}_{k}(y))=L_{k}(y_{0,k}+d_{y_{0,k}}N_{h_{k}}^{-1}y),$
has uniform $C^{3,\gamma}$ bounds on $B_{k}(0)$ and satisfies $|D^{3}\widetilde{h}_{k}(0)|=1$. By passing to a subsequence we may assume that $y_{0,k}\to y_{0,\infty}\in B$. Combining with bounds for $a_{k,ij}$, $b_{k}$, $L_{k}$ as well as a standard diagonal argument shows that by taking a subsequence we can extract a convex limit $\widetilde{h}_{\infty}\colon\mathbb{R}^{n}\to\mathbb{R}$ in $C_{\operatorname{loc}}^{3,\gamma/2}$, satisfying $|D^{3}\widetilde{h}_{\infty}(0)|=1$ and
$\widetilde{a}_{\infty,ij}\widetilde{h}_{\infty,ij}(y)+\widetilde{b}_{\infty}\det(D^{2}\widetilde{h}_{\infty}(y))=L_{\infty}(y_{0,\infty})$
with constant coefficients $\widetilde{a}_{\infty,ij}$, $\widetilde{b}_{\infty}$ due to (5.15). Since $L_{\infty}(y_{0,\infty})\geqslant\delta>0$, after a linear change of coordinates and scaling of $\widetilde{h}_{\infty}$ we can apply the Liouville rigidity result [CS17, Lemma 31] to conclude that $\widetilde{h}_{\infty}$ is a quadratic polynomial. This contradicts the fact that $|D^{3}\widetilde{h}_{\infty}(0)|=1$, completing the proof. ∎
6. A numerical criterion for extremal Kähler metrics
In this section, we give a proof of Theorem 1.7 by adapting Theorem 1.6 to the arguments in [DJ25] [^3] More precisely, Delcroix–Jubert [DJ25] obtain an existence result for $(\mathrm{v},\mathrm{w})$-weighted cscK metrics in their context, and extremal Kähler metrics arises as the special case with $\mathrm{v}=1$ and $\mathrm{w}$ being the affine linear function correspoinding to the extremal vector field $v_{\rm ext}$.. First, let us consider a general compact complex manifold $X$ and a maximal compact torus $T\subset\operatorname{Aut}_{\rm red}(X)$. Let $\widehat{\chi}$ be a $T$-invariant Kähler form, and let $\beta$ denote its Kähler class. We choose the extremal vector field $v_{\rm ext}$ so that $\mathrm{Im}(v_{\rm ext})\in\mathfrak{t}$. The following discussion is entirely based on [DJ25]. However, since there are differences in conventions such as the definitions of functionals, we briefly summarize and unify them in our setting for the reader’s convenience. We define the modified $K$-energy
$K_{v_{\rm ext}}(\phi)=-\int_{0}^{1}\int_{X}\phi\big(S(\chi_{t\phi})-\overline{S}-\theta_{v_{\rm ext}}^{X}(\chi_{t\phi})\big)\chi_{t\phi}^{n}dt.$
where $S(\chi_{t\phi})$ denotes the scalar curvature of $\chi_{t\phi}$, and $\overline{S}=\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}$ its average. A critical point of the functional $K_{v_{\rm ext}}$, i.e. a solution $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ to the equation
$S(\chi_{\phi})-\overline{S}=\theta_{v_{\rm ext}}^{X}(\chi_{\phi})$
is called an extremal Kähler metric (cf. [Cal82]). By the definition of $v_{\rm ext}$ the modified $K$-energy $K_{v_{\rm ext}}$ is invariant under the action of the complexified torus $T^{\mathbb{C}}$. Let $v$ be a holomorphic vector field on $X$ such that $\mathrm{Im}(v)\in\mathfrak{t}$, and let $\rho$ be a real closed $(1,1)$-form on $X$, which is not necessarily Kähler. For later use, it is convenient to extend (2.7) to $\rho$ as well:
$J_{v}^{\rho}(\phi)=\int_{0}^{1}\int_{X}\phi\big(\operatorname{Tr}_{\chi_{t\phi}}\rho-c_{[\rho]}-\theta_{v}^{X}(\chi_{t\phi})\big)\chi_{t\phi}^{n}dt,\quad c_{[\rho]}:=\frac{n[\rho]\cdot\beta^{n-1}}{\beta^{n}}.$
It is easy to verify that $J_{s_{1}v_{1}+s_{2}v_{2}}^{s_{1}\rho_{1}+s_{2}\rho_{2}}=s_{1}J_{v_{1}}^{\rho_{1}}+s_{2}J_{v_{2}}^{\rho_{2}}$ for all $s_{1},s_{2}\in\mathbb{R}$. We set
$J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi):=\inf_{\tau\in T^{\mathbb{C}}}J^{\widehat{\chi}}(\phi_{\tau}),$
where the potential $\phi_{\tau}\in\mathcal{H}(X,\widehat{\chi})^{T}$ satisfies $\tau^{\ast}\chi_{\phi}=\chi_{\phi_{\tau}}$. By Remark 2.18, the functional $J^{\widehat{\chi}}$ is comparable with $I$, $J$ and hence $d_{1}$-distance as well. In particular, for a fixed $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, there exists $\tau\in T^{\mathbb{C}}$ that attains the infimum in the definition of $J_{T^{\mathbb{C}}}^{\widehat{\chi}}$ [DJ25, Theorem 3.7]. We say that the modified $K$-energy $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive if there exist constants $C_{1},C_{2}>0$ such that
$K_{v_{\rm ext}}(\phi)\geqslant C_{1}J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)-C_{2}$
for all $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$. We define the relative entropy $H$ by
$H(\phi):=\int_{X}\log\frac{\chi_{\phi}^{n}}{\widehat{\chi}^{n}}\chi_{\phi}^{n}.$
Using these functionals, the modified $K$-energy $K_{v_{\rm ext}}$ can be expressed as
$K_{v_{\rm ext}}(\phi)=H(\phi)+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}(\phi).$ (6.1)
This is well-known as the Chen–Tian formula (cf. [Che00, Tia97a, Tia97b]). Define the $T^{\mathbb{C}}$-reduced delta invariant of the Kähler class $\beta$ by
$\delta_{\beta}^{T^{\mathbb{C}}}:=\sup\bigg\{\delta\in\mathbb{R}\bigg|\inf_{\phi\in\mathcal{H}(X,\widehat{\chi})^{T}}\big(H(\phi)-\delta J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)\big)\in\mathbb{R}\bigg\}.$ (6.2)
Note that the value $\delta_{\beta}^{T^{\mathbb{C}}}$ does not depend on the choice of reference form $\widehat{\chi}\in\beta$, as can be verified by an argument similar to the proof of Theorem 1.3. By using Tian’s alpha invariant, one can also see that $\delta_{\beta}^{T^{\mathbb{C}}}>0$ [DJ25, Section 3.2]. In view of (6.1), the invariant $\delta_{\beta}^{T^{\mathbb{C}}}$ quantitatively measures the coercivity of the modified $K$-energy $K_{v_{\rm ext}}$. Let $\varepsilon>0$ be a constant. For each $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, let us take $\tau\in T^{\mathbb{C}}$ that minimizes $J^{\widehat{\chi}}(\phi_{\tau})$ so that $J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi_{\tau})=J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)=J^{\widehat{\chi}}(\phi_{\tau})$. As in [DJ25, Lemma 5.4], using the $T^{\mathbb{C}}$-invariance of $K_{v_{\rm ext}}$, Chen–Tian formula (6.1), and definition of $\delta_{\beta}^{T^{\mathbb{C}}}$, we obtain
$K_{v_{\rm ext}}(\phi)$ $=K_{v_{\rm ext}}(\phi_{\tau})$
$=H(\phi_{\tau})+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}(\phi_{\tau})$
$\geqslant\bigg(\frac{\varepsilon}{2}+\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon\bigg)J^{\widehat{\chi}}(\phi_{\tau})+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}(\phi_{\tau})-C$
$=\frac{\varepsilon}{2}J^{\widehat{\chi}}(\phi_{\tau})+J^{(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\widehat{\chi}}(\phi_{\tau})+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}(\phi_{\tau})-C$
$=\frac{\varepsilon}{2}J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)+J_{-v_{\rm ext}}^{(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\widehat{\chi}-\operatorname{Ric}(\widehat{\chi})}(\phi_{\tau})-C.$
Thus, showing that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive reduces to the problem of establishing a lower bound for the functional $J_{-v_{\rm ext}}^{(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\widehat{\chi}-\operatorname{Ric}(\widehat{\chi})}$. Moreover, if the class $\alpha:=(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler, then we can choose a Kähler form $\omega\in\alpha$, and define a modified $J$-functional
$J_{-v_{\rm ext}}^{\omega}(\phi)=\int_{0}^{1}\int_{X}\phi\big(\operatorname{Tr}_{\chi_{t\phi}}\omega-n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)+\overline{S}-\theta_{-v_{\rm ext}}^{X}(\chi_{t\phi})\big)\chi_{t\phi}^{n}dt.$ (6.3)
[…]
$\operatorname{Tr}_{\chi_{\phi}}\omega=n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\overline{S}+\theta_{-v_{\rm ext}}^{X}(\chi_{\phi}).$ (6.4)
Therefore, by applying Theorem 1.6 to the equation (6.4) in the toric case, we know that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive under the assumptions in Theorem 1.7. Then it follows from [NJL25, Theorem 1.2] that the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric. This completes the proof of Theorem 1.7.

### v4

@@
 In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16].
-We show that the solvability of the modified $J$-equation is equivalent to the coercivity of the modified $J$-functional on compact Kähler manifolds.
-For smooth projective toric varieties we establish a Nakai–Moishezon type criterion for the existence of solutions, extending the results of Collins–Székelyhidi [CS17].
-As a possible application, by combining our result with Delcroix–Jubert [DJ25], we further discuss a numerical sufficient condition for the existence of extremal Kähler metrics on smooth projective toric varieties.
+We first show that, on compact Kähler manifolds, the solvability of the modified $J$-equation is equivalent to the coercivity of the modified $J$-functional.
+Motivated by this characterization, we formulate a Nakai–Moishezon type criterion for the existence of solutions to the modified $J$-equation on general compact Kähler manifolds.
+We then verify this conjectural criterion in the case of smooth projective toric varieties.
+This extends the work of Collins–Székelyhidi [CS17] and provides further evidence for the expected algebro-geometric nature of the modified $J$-equation.
+As a potential application, we combine our results with Delcroix–Jubert [DJ25].
+Assuming our conjectural Nakai–Moishezon type criterion holds in general, we obtain a numerical sufficient condition for the existence of extremal Kähler metrics on arbitrary compact Kähler manifolds.
 Key words and phrases:
@@
 $i_{v}\chi_{\phi}=\frac{\sqrt{-1}}{2\pi}\bar{\partial}\theta_{v}^{X}(\chi_{\phi}),\quad\int_{X}\theta_{v}^{X}(\chi_{\phi})\chi_{\phi}^{n}=0.$ (1.1)
+1.1. Background and motivation
 Finding canonical metrics has been one of the central themes in Kähler geometry.
-Especially several equations involving Hamiltonians have been studied extensively in the last few decades (e.g.
+Equations such as the constant scalar curvature Kähler (cscK) equation, the Kähler–Einstein equation, and their various twisted or weighted analogues reveal deep interactions between complex geometry, algebraic stability, and variational principles (e.g.
 [BJ25, FO19, Lah19, Li21, Mab03, TZ00, Zhu00]).
-Hamiltonians of holomorphic vector fields encode the symmetries of the manifold with respect to a given Kähler metric.
-When studying canonical metrics, such as Kähler–Einstein or constant scalar curvature Kähler (cscK) metrics, it is natural to take these Hamiltonians into account, since the existence and uniqueness of such metrics are often obstructed by holomorphic automorphisms.
-By introducing Hamiltonians of holomorphic vector fields, one can treat the variational problem even without a canonical metric.
-In this paper, we study the modified $J$-equation introduced by Li–Shi [LS16]:
+Among these, the $J$-equation, introduced by Donaldson [Don03] and Chen [Che00], plays a crucial role in understanding the relation between energy properness and the existence of cscK metrics.
+It also serves as an important model in the continuity approach to cscK metrics [CC18, CC21a, CC21b].
+An important progress in the analysis of the $J$-equation was achieved by Song–Weinkove [SW08] who proved that the existence of a subsolution is equivalent to the solvability of the equation.
+Building on this insight, Collins–Székelyhidi [CS17] established a fundamental correspondence between the solvability of the $J$-equation and the coercivity of the associated functional, providing a model for the continuity method in Kähler geometry.
+A further breakthrough was made by G.
+Chen [Che21], who established a connection between the solvability of the $J$-equation and algebro-geometric stability.
+Using Chen’s result [Che21], Song [Son20] obtained a Nakai–Moishezon type numerical criterion for the solvability of the $J$-equation, confirming the Lejmi–Székelyhidi conjecture [LS15].
+However, the usual $J$-equation does not incorporate the effect of holomorphic automorphisms, which often obstruct the existence or uniqueness of canonical metrics.
+To address this, Li–Shi [LS16] introduced the modified $J$-equation :
 $\omega\wedge\chi_{\phi}^{n-1}=\frac{1}{n}(c_{X}+\theta_{v}^{X}(\chi_{\phi}))\chi_{\phi}^{n},$ (1.2)
@@
 is independent of the choice of $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, and the existence of a solution to (1.2) easily leads to $m_{X}>0$ (see Section 2.2).
-In particular, if $T$ is trivial then the equation (1.2) matches up with the $J$-equation introduced by Donaldson [Don03] and Chen [Che00].
-Although the $J$-equation itself is an intriguing equation, it has sometimes been studied in relation to cscK metrics, as it provides the starting point of the continuity method for solving the cscK equation [CC18, CC21a, CC21b].
-In the modified case, Li–Shi [LS16] showed that such a solution $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$ to (1.2) exists if and only if there exists $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$ such that
-$\big(c_{X}+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0,$ (1.5)
-where the positivity of (n1,n1)(n-1,n-1)-forms is understood in the sense of [Dem12, Chapter III, Section 1.A].
-The solutions to the equation (1.2) are characterized as the critical points of a strictly convex function $J_{v}^{\omega}$ on $\mathcal{H}(X,\widehat{\chi})^{T}$, called the modified $J$-functional.
-On the other hand, the modified $K$-energy $K_{v_{\rm ext}}$ is a convex functional on the space of Kähler potentials whose critical points are extremal Kähler metrics.
-Motivated by [LS16], Delcroix–Jubert [DJ25] recently considered a modified version of the Chen–Tian formula, relating these functionals, and proved that the existence of a solution to the modified $J$-equation implies the $T^{\mathbb{C}}$-coercivity of the modified $K$-energy.
+They proved that the existence of a subsolution $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$:
+$\big(c_{X}+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0$ (1.5)
+is equivalent to the solvability of the equation (1.2). This equivalence was obtained by studying the modified $J$-flow
+$\frac{d}{dt}\phi_{t}=-\operatorname{Tr}_{\chi_{\phi_{t}}}\omega+c_{X}+\theta_{v}^{X}(\chi_{\phi_{t}})$ (1.6)
+with initial data $\phi_{0}\in\mathcal{H}(X,\widehat{\chi})^{T}$.
+The modified $J$-equation is closely related to the study of extremal Kähler metrics, since the latter can be viewed as critical points of the modified $K$-energy.
+In this sense, the modified $J$-theory may serve as an analytic model for understanding extremal and, more generally, weighted cscK metrics (see Section 1.3).
+1.2. Main results
+The purpose of this paper is to establish the analytic and numerical foundations of the modified $J$-theory, and to clarify how it extends the previous framework developed by [CS17, LS16].
 For later arguments, it is convenient to extend (1.2) to a more general equation of the form:
@@
 $b=\frac{c\beta^{n}-n\alpha\cdot\beta^{n-1}}{\alpha^{n}}.$ (1.8)
-In most cases, we assume that $b\geqslant 0$, but sometimes allow $b$ to be slightly negative. First, we extends the result [LS16, Theorem 3.3] to the generalized equation (1.6) as follows:
+In most cases, we assume that $b\geqslant 0$, but sometimes allow $b$ to be slightly negative. We extends the result [LS16, Theorem 3.3] to the generalized equation (1.7) as follows:
 Theorem 1.1.
@@
 $\big(c+\theta_{v}^{X}(\chi_{\underline{\phi}})\big)\chi_{\underline{\phi}}^{n-1}-(n-1)\omega\wedge\chi_{\underline{\phi}}^{n-2}>0.$ (1.9)
-In the study of canonical Kähler metrics, a fundamental question is whether the solvability of an equation can be related to the properness or coercivity of a certain energy functional.
-For instance, Tian’s properness conjecture [Tia94] predicted that the existence of Kähler–Einstein metrics is equivalent to the properness of the $K$-energy.
-When the manifold has trivial automorphism group, Ding–Tian [DT92] proved one direction (properness $\Rightarrow$ existence), and Tian [Tia97a] confirms a converse direction (existence $\Rightarrow$ properness).
-In general case, Darvas–Rubinstein [DR17] proposed a refined version of Tian’s properness conjecture and proved it by exploiting a general framework in Finsler geometry.
-For constant scalar curvature Kähler metrics, a version of Tian’s properness conjecture was proved by Berman–Darvas–Lu [BDL20] and Chen–Cheng [CC18, CC21a, CC21b].
-For complex Hessian equations, Collins–Székelyhidi [CS17] obtained a similar result for the $J$-equation, and Chu–Lee [CL21] for the hypercritical deformed Hermitian–Yang–Mills equation.
-Motivated by above works, we show the equivalence between the existence and coercivity for the modified $J$-equation:
+Our first main result is an analogue of the existence-coercivity correspondence proved for the usual $J$-equation in [CS17].
 Theorem 1.2.
@@
 For the precise definition of $J_{v}^{\omega}$ and coercivity, see Section 2.4.
-To prove the above theorem, we use the local smoothing and gluing method established in [Che21, CS17].
-A new issue arising in the modified case is that even if we approximate the holomorphic vector field $v$ locally by vector fields with constant coefficients, to deal with the approximation error terms, we need a uniform gradient estimate for the potentials along the modified $J$-flow.
-To avoid this, we consider the product of $X$ with an annulus, add a rotational vector field $w$ to $v$ so as to eliminate the zeroes.
-We take holomorphic coordinates on which $v+w$ has constant coefficients by using the rectification theorem.
-Then we perform smoothing, gluing and finally averaging in the rotational direction to cancel out the contribution of $w$ (see Section 4.1.1 for more details).
+This result shows that the coercivity of the modified $J$-functional completely characterizes solvability, in the same spirit as the coercivity criterion for various canonical metrics (e.g.
+[BDL20, CC18, CC21a, CC21b, CL21, DR17, DT92, Tia97a]).
 In practice, it seems to be difficult to produce $\underline{\phi}\in\mathcal{H}(X,\widehat{\chi})^{T}$ satisfying (1.5).
@@
 where $d^{c}:=\sqrt{-1}(\bar{\partial}-\partial)$. So the desired statement follows from the equivariant Chern–Weil theory (cf. [BGV92]).
-[…]
+Remark 1.6.
+In particular, when $T$ is trivial and $n=2$, the modified $J$-equation (1.2) reduces to a complex Monge-Ampère equation:
+$(c_{X}\chi_{\phi}-\omega)^{2}=\omega^{2}.$
+Consequently, Conjecture 1.4 follows directly from the criterion of Demailly–Păun [DP04] and [Yau78].
+In contrast, when $T$ is non-trivial, even for surfaces ($n=2$), the modified $J$-equation (1.2) does not reduce to a complex Monge-Ampère equation, and Conjecture 1.4 remains genuinely nontrivial.
+In the case where $T$ is trivial, this was proposed by Lejmi–Székelyhidi [LS15] and subsequently established by Song [Son20], using [Che21] as a key ingredient.
+We also note that, when the condition (1.10) fails, the formation of singularities of the modified $J$-flow (1.6) in the Calabi ansatz has been studied by Sivaram [Siv24].
+Focusing on the toric setting allows for more explicit results.
+Theorem 1.7 confirms Conjecture 1.4 in this case, providing a concrete demonstration of the expected behavior under these assumptions.
 Theorem 1.7.
@@
 where $\chi\in\beta$ is any $T$-invariant Kähler form.
-When dealing with the contribution of the holomorphic vector field $v$, one need to handle carefully the fact that the normalization of the Hamiltonian (1.1) differs for each subvariety $Y$.
-Also, when gluing subsolutions, we need to control the derivatives of potential functions along $v$ near each $D_{i}$, using a $T$-equivariant linearization of $\mathcal{O}_{X}(D_{i})$.
+As in [CS17, Theorem 3], the proof proceeds by analyzing the behavior outside the union of toric divisors $D=\cup_{i}D_{i}$, adapting their method to the present context.
+Unlike [CS17], we need to control the Hamiltonian near each $D_{i}$ during the gluing process, by using a $T$-equivariant linearization of $\mathcal{O}_{X}(D_{i})$.
 For this reason, we make the additional assumption that $X$ is projective.
-As an application, we also obtain:
-Theorem 1.7.
-Let $X$ be an $n$-dimensional smooth projective toric variety with real torus $T$, and let $\widehat{\chi}$ be a $T$-invariant Kähler form on $X$ with Kähler class $\beta$.
-Let $\delta_{\beta}^{T^{\mathbb{C}}}$ be the $T^{\mathbb{C}}$-reduced delta invariant of $\beta$ defined by (6.2).
-Let $v_{\rm ext}$ be the extremal vector field associated with $\beta$ such that $\mathrm{Im}(v_{\rm ext})\in\mathfrak{t}$.
-Assume that there exists a constant $\varepsilon>0$ satisfying the following properties:
-  1. (1)
-    The class $(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler.
-  2. (2)
-    For any $T$-invariant Kähler forms $\chi\in\beta$,
-    $\max_{X}\theta_{v_{\rm ext}}^{X}(\chi)<n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}.$
-  3. (3)
-    For any $p$-dimensional irreducible toric subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ and any $T$-invariant Kähler forms $\chi\in\beta$,
-    $\int_{Y}\bigg[\bigg((n-p)(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}-\theta_{v_{\rm ext}}^{X}(\chi)\bigg)\chi^{p}+p\chi^{p-1}\cdot c_{1}(X)\bigg]>0.$
-Then the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric.
-Remark 1.8.
-We emphasize that the above theorem is essentially established in [DJ25, Theorem 1.4].
-Our contribution is limited in the toric case.
-Specifically, in the course of the proof, we need to show the existence of a solution to a certain modified $J$-equation, which we reformulate in numerical terms by using Theorem 1.6.
-Thus, if Conjecture 1.4 is true, Theorem 1.7 still holds for arbitrary compact Kähler manifolds.
-Also, it should be noted that in the toric case, a stronger statement holds: the conclusions remain valid even if $\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon$ in assumptions (1)–(3) is replaced by $\delta_{\beta}^{T^{\mathbb{C}}}$.
-This is due to the fact that, in (3), the object under considerations are restricted to toric subvarieties, which form a finite set.
-We note that in the case $v=0$, Theorem 1.2, Theorem 1.3 and Theorem 1.6 were previously established in [CS17]. Also, a variant of Theorem 1.7 was obtained in [Che21, Corollary 1.5].
+1.3. Relation to extremal Kähler metrics
+Recall that extremal Kähler metrics are precisely the critical points of the modified $K$-energy $K_{v_{\rm ext}}$, which can be decomposed into the relative entropy and the modified $J$-functional.
+This decomposition is exactly the Chen–Tian formula [Che00, Tia97a, Tia97b]:
+$K_{v_{\rm ext}}=H+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}$.
+Using this formula, Li–Shi [LS16] established a sufficient condition related to the modified $J$-equation under which the modified $K$-energy is coercive.
+Building on this perspective, Delcroix–Jubert [DJ25] further showed that the $T^{\mathbb{C}}$-coercivity of the modified $K$-energy can be reduced to the existence problem for the modified $J$-equation.
+This reduction is a key step in connecting extremal Kähler metrics with solutions of the modified $J$-equation.
+Their work, however, addresses a more general setting, namely so-called weighted cscK metrics, which have recently attracted considerable attention in Kähler geometry (e.g.
+[AJL23, BJ25, HL24, Lah19, NJL25]).
+It is natural to expect that the theory of the modified $J$-equation can be further adapted in accordance with their framework, potentially yielding applications to the existence problem for such general canonical metrics.
+We leave a systematic exploration of such general cases to future work.
+At the same time, the combination of our results with [DJ25] provides a framework for potential applications to extremal Kähler metrics.
+We explore these applications and related discussions in Section 6.
+1.4. Contributions
+The main contributions of this paper concern the theory of the modified $J$-equation, with the principal new analytic ingredients being the local smoothing and gluing.
+Our analysis builds on the techniques developed in [Che21, CS17], which we further adapt to handle the lack of $T$-invariance and the degeneracy of the automorphism flow generated by the holomorphic vector field $v$.
+These issues form a major obstacle to the construction of subsolutions through local smoothing and gluing procedures.
+To overcome them, we employ several analytic tools.
+The first contribution consists of two techniques: the introduction of averaging with respect to the torus action, which restores $T$-invariance, and a careful control of the Hamiltonian near $T$-invariant subvarieties.
+These together make it possible to carry out the gluing procedures in a way compatible with the $T$-action (cf.
+Section 2.3 and Theorem 5.2).
+The second is the annulus trick, used in the construction of solutions in Theorem 1.2.
+When establishing the local smoothing estimate, the key is to work in local coordinates that are compatible with the automorphism flow generated by the holomorphic vector field $v$, so that the modified structure is preserved.
+However, such coordinates can not be defined consistently near the zero set of $v$, where the automorphism action degenerates.
+To avoid this, we consider the product space $X\times A$, where $A=\{\tau\in\mathbb{C}\mid 1/2<|\tau|<2\}$ is an annulus equipped with a rotational vector field $w=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$.
+By lifting $v$ to $X\times A$ and adding $w$, we obtain a flow without fixed points.
+This extended flow allows us to define local coordinates adapted to it, which are then used to perform smoothing and gluing constructions.
+Combined with an averaging argument along the rotational direction, this yields a subsolution on the slice $X\simeq X\times\{1\}\subset X\times A$.
+This trick is not merely a computational convenience; it is essential in that it allows us to bypass the highly nontrivial analytic problem of establishing uniform gradient estimates along the modified $J$-flow (see Section 4.1.1 for more details).
+1.5. Organization of the paper
 The organization of this paper is as follows.
-In Section 2 we collect some fundamental properties that will be used throughout this paper.
-In Section 3 we establish the long-time existence of the modified $J$-flow.
-In particular we prove the convergence of the flow when $b\geqslant 0$ under the existence of solusolutions, which gives a proof of Theorem 1.1.
-In Section 4 we show that the existence of solutions is equivalent to the coercivity of the functional $J_{v}^{\omega}$ (Theorem 1.2).
-As a corollary, we also obtain a proof of Theorem 1.3.
-In Section 5 we establish a Nakai–Moishezon type criterion for smooth projective toric varieties, which provides the proof of Theorem 1.6.
-Finally, in Section 6, we give the proof of Theorem 1.7.
+In Section 2 we review several basic properties that will be used throughout the paper.
+Section 3 is devoted to the study of the long-time behavior of the modified $J$-flow; in particular, we prove its convergence when $b\geqslant 0$ under the existence of subsolutions, thereby establishing Theorem 1.1.
+In Section 4, we show that the existence of solutions is equivalent to the coercivity of the functional $J_{v}^{\omega}$ (Theorem 1.2), from which Theorem 1.3 follows as a corollary.
+Section 5 develops a Nakai–Moishezon type criterion for smooth projective toric varieties, providing the proof of Theorem 1.7.
+Finally, in Section 6, we discuss how Conjecture 1.4 can be combined with the results of Delcroix–Jubert [DJ25] to address the existence problem for extremal Kähler metrics.
 Acknowledgment.
@@
 In this subsection we study general properties for subsolutions (1.5) related to the regularized maximum and averaging over the torus $T$.
-We first discuss how the subsolution coindition can be expressed in terms of the operator $P_{\omega}$.
+We first discuss how the subsolution condition can be expressed in terms of the operator $P_{\omega}$.
 Proposition 2.11.
@@
 By the convexity of $M_{\eta}$, (2.4) and (2.5) we observe that
-$\chi\geqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j})>0.$
+$\chi\geqslant\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot\bigg(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\bigg)>0.$
 The term $\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}\cdot\big(\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}\big)$ is a weighted average of $\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\varphi_{j}$ with $\sum_{j}\frac{\partial M_{\eta}}{\partial t_{j}}=1$.
@@
 Since $\omega$ is $T$-invariant, the Kähler form $\tau^{\ast}\chi$ can be identified with a positive definite Hermitian matrix by using an $\omega$-orthonormal basis at each point on $X$.
-So the convexity of $P_{\omega}$ implies that
+So the convexity of the function $P$ implies that
 $P_{\omega}(\widetilde{\chi})\leqslant\int_{\tau\in T}\tau^{\ast}P_{\omega}(\chi)dT.$
@@
 $0<-\partial_{pp}F_{-\varepsilon}(A_{t})=-\partial_{p}f_{-\varepsilon}(\lambda)=\frac{1}{\lambda_{p}^{2}}-\frac{\varepsilon}{\lambda_{p}S_{n}(\lambda)}.$
-Since $\lambda_{i}$ is uniformly bounded from blow this shows that the right hand side is bounded from above by some uniform constant $C_{4}>0$.
+Since $\lambda_{i}$ is uniformly bounded from below this shows that the right hand side is bounded from above by some uniform constant $C_{4}>0$.
 Therefore if we set $G:=\log\operatorname{Tr}_{\omega}\chi_{t}-2C_{4}t$ then for any $T\in(0,T_{\max})$ and maximum point $(x_{0},t_{0})$ of $G$ on $X\times[0,T]$ with $t_{0}>0$ we must have
@@
 $I(u_{k})\leqslant\frac{J_{v,-\varepsilon}^{\omega}(u_{k})+B_{c}}{\delta_{c}/2}\leqslant\frac{J_{v,-\varepsilon}^{\omega}(0)+B_{c}}{\delta_{c}/2}\leqslant C$ (4.3)
-[…]
+for some uniform constant $C>0$ (independent of $\varepsilon$ and $k$).
+Hence, by choosing a subsequence, we may assume that $u_{k}\to u_{\infty}\in\mathcal{E}^{1}(X,\widehat{\chi})^{T}$ in $L^{1}$ and $\chi_{u_{k}}\to\chi_{u_{\infty}}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{\infty}$ weakly as $k\to\infty$, where $\mathcal{E}^{1}(X,\widehat{\chi})^{T}$ denotes the space of $T$-invariant $\widehat{\chi}-$PSH functions with finite energy (see Remark 2.18 and [BBGZ13, Lemma 3.3]).
+So by [GZ07, Corollary 1.8] the function $u_{\infty}$ has zero Lelong numbers.
 4.1.1. Local smoothing
 The Kähler current $\chi_{u_{\infty}}$ can be regarded as a solution to (1.2) in a weak sense.
-Now we establish estimates for local smoothing of $\chi_{u_{\infty}}$ in Euclidean balls.
-The main difference from the usual $J$-equation lies in the smoothing estimates for the Hamiltonian, which are not straightforward.
+Now we establish estimates for local smoothing of $\chi_{u_{\infty}}$.
+Following the general strategy in [Che21, CS17], we locally approximate the Kähler form $\omega$ and the holomorphic vector field $v$ by constant-coefficient ones in each coordinate chart.
+This approximation makes the smoothing operator commute with differentiation.
+For the Hessian term, the desired estimate then follows from the monotonicity and convexity of the function $P$.
+However, the argument for the Hamiltonian term is more delicate.
 Even if $v$ is locally approximated by a constant-coefficient vector field, controlling the resulting error term requires a uniform gradient bound $|\nabla u_{k}|_{g}\leqslant C$.
+Obtaining this bound (without assuming the existence of a subsolution) seems to be a highly challenging problem in its own right.
 While this error term could be eliminated using the rectification theorem, this approach can not be applied directly since $v$ must have a zero on $X$.
-To address this issue, one may consider a product manifold $\mathcal{X}:=X\times\mathbb{C}$ instead of $X$: let $\omega_{\mathbb{C}}:=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}|\tau|^{2}$ be the standard Kähler form on $\mathbb{C}$ and $w:=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$ a holomorphic vector field on $\mathbb{C}$ whose real part generates the action of the real torus $T^{\prime}\simeq\mathbb{S}^{1}$ on $\mathbb{C}$ by rotations.
+To address this issue, we perform the annulus trick as follows: let us consider a product manifold $\mathcal{X}:=X\times\mathbb{C}$.
+Let $\omega_{\mathbb{C}}:=\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}|\tau|^{2}$ be the standard Kähler form on $\mathbb{C}$ and $w:=\sqrt{-1}\tau\frac{\partial}{\partial\tau}$ a holomorphic vector field on $\mathbb{C}$ whose real part generates the action of the real torus $T^{\prime}\simeq\mathbb{S}^{1}$ on $\mathbb{C}$ by rotations.
 Set
@@
 The holomorphic vector field $v+w$ still has constant coefficients in $(z^{1},\ldots,z^{n+1})$, which we denote by $c(v+w)\in\mathbb{C}^{n+1}$.
-Also a simple computation shows that $|c(v+w)|=|v+w|_{g_{\mathcal{X}}}(p)\leqslant\sup_{\mathcal{X}_{3}}|v+w|_{\omega_{\mathcal{X}}}$.
+Also a simple computation shows that $|c(v+w)|=|v+w|_{g_{\mathcal{X}}}(p)\leqslant\sup_{\mathcal{X}_{3}}|v+w|_{g_{\mathcal{X}}}$.
 Taking this into the account, we can choose sufficiently small constants $R>0$, $\sigma>0$ and a finite covering $\{B_{j,4R}\}_{j\in\mathcal{J}}$ of $\mathcal{X}_{2}$ satisfying the following properties:
@@
 Lemma 4.2.
-For any $\varepsilon\in(0,\varepsilon_{7})$ we have
-$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\leqslant F_{\omega,-\varepsilon}(\chi_{u_{k}})+\varepsilon^{2},$
-where the quantity in the left-hand side is calculated on $\mathcal{X}$.
+For any $\varepsilon\in(0,\varepsilon_{7})$, we have
+$R_{\omega_{\mathcal{X}}}\bigg(\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}u_{k,j}\bigg)\leqslant F_{\omega,-\varepsilon}(\chi_{u_{k}})+\varepsilon^{2}$
+on $B_{j,4R}$ for all $j\in\mathcal{J}$, where the quantity in the left-hand side is calculated on $\mathcal{X}$.
 Proof.
@@
 $\varphi_{j,r}(z)\leqslant\varphi_{i,r}(z)-\frac{\varepsilon^{2}R^{2}}{2}.$
-The above discussion prepares the ground for constructing a subsolution on $X$. We are now ready to begin the proof.
+The above discussion prepares the ground for constructing a subsolution on $X$. We now proceed to the proof.
 Proof of “coercivity $\Rightarrow$ existence” in Theorem 1.2.
@@
 $P_{\omega}(\chi)-\theta_{v}^{X}(\widehat{\chi})-\mathrm{Re}(v)(\widetilde{\varphi}|_{X})<c_{X},$
-where we used Proposition 2.2 and the fact that $\mathrm{Re}(v)$ is tangent to $X$. Hence, by Proposition 2.15, if we set
+where we used Proposition 2.2 (2) and the fact that $\mathrm{Re}(v)$ is tangent to $X$. Hence, by Proposition 2.15, if we set
 $\phi:=\int_{\tau\in T}\tau^{\ast}\widetilde{\varphi}|_{X}dT,\quad\widetilde{\chi}:=\widehat{\chi}+\frac{\sqrt{-1}}{2\pi}\partial\bar{\partial}\phi,$
@@
 for some constant $\delta>0$.
-Adding an affine linear function to $g$ amounts to a translation of the moment polytope $P$.
+Adding an affine linear function to $g$ amounts to a parallel translation of the moment polytope $P$.
 Since the equation (5.7) and the condition (5.8) remain unchanged under this operation, we may assume that $g(0)=0$ and $\nabla g(0)=0$.
@@
 with constant coefficients $\widetilde{a}_{\infty,ij}$, $\widetilde{b}_{\infty}$ due to (5.15).
-Since $L_{\infty}(y_{0,\infty})\geqslant\delta>0$, after a linear change of coordinates and scaling of $\widetilde{h}_{\infty}$ we can apply the Liouville rigidity result [CS17, Lemma 31] to conclude that $\widetilde{h}_{\infty}$ is a quadratic polynomial.
+Since $L_{\infty}(y_{0,\infty})\geqslant\delta>0$, after a linear change of coordinates, we can apply the Liouville rigidity result [CS17, Lemma 31] to conclude that $\widetilde{h}_{\infty}$ is a quadratic polynomial.
 This contradicts the fact that $|D^{3}\widetilde{h}_{\infty}(0)|=1$, completing the proof.
@@
 6. A numerical criterion for extremal Kähler metrics
-In this section, we give a proof of Theorem 1.7 by adapting Theorem 1.6 to the arguments in [DJ25] [^3] More precisely, Delcroix–Jubert [DJ25] obtain an existence result for $(\mathrm{v},\mathrm{w})$-weighted cscK metrics in their context, and extremal Kähler metrics arises as the special case with $\mathrm{v}=1$ and $\mathrm{w}$ being the affine linear function correspoinding to the extremal vector field $v_{\rm ext}$..
-First, let us consider a general compact complex manifold $X$ and a maximal compact torus $T\subset\operatorname{Aut}_{\rm red}(X)$.
+Let $X$ be an $n$-dimensional compact Kähler manifold and $T\subset\operatorname{Aut}_{\rm red}(X)$ a maximal compact torus.
 Let $\widehat{\chi}$ be a $T$-invariant Kähler form, and let $\beta$ denote its Kähler class.
@@
 $K_{v_{\rm ext}}(\phi)=H(\phi)+J_{-v_{\rm ext}}^{-\operatorname{Ric}(\widehat{\chi})}(\phi).$ (6.1)
-This is well-known as the Chen–Tian formula (cf. [Che00, Tia97a, Tia97b]). Define the $T^{\mathbb{C}}$-reduced delta invariant of the Kähler class $\beta$ by
+This is well-known as the Chen–Tian formula [Che00, Tia97a, Tia97b]. Define the $T^{\mathbb{C}}$-reduced delta invariant of the Kähler class $\beta$ by
 $\delta_{\beta}^{T^{\mathbb{C}}}:=\sup\bigg\{\delta\in\mathbb{R}\bigg|\inf_{\phi\in\mathcal{H}(X,\widehat{\chi})^{T}}\big(H(\phi)-\delta J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)\big)\in\mathbb{R}\bigg\}.$ (6.2)
@@
 In view of (6.1), the invariant $\delta_{\beta}^{T^{\mathbb{C}}}$ quantitatively measures the coercivity of the modified $K$-energy $K_{v_{\rm ext}}$.
-Let $\varepsilon>0$ be a constant.
+As an application to the existence problem of extremal Kähler metrics, we now prove the following theorem:
+Theorem 6.1.
+Let $X$ be an $n$-dimensional compact Kähler manifold, and $T$, $v_{\rm ext}$, $\widehat{\chi}$, $\beta$ as above.
+Assume that Conjecture 1.4 holds, and that there exists a constant $\varepsilon>0$ satisfying the following properties:
+  1. (1)
+    The class $(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler.
+  2. (2)
+    For any $T$-invariant Kähler forms $\chi\in\beta$,
+    $\max_{X}\theta_{v_{\rm ext}}^{X}(\chi)<n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}.$
+  3. (3)
+    For any $T$-invariant $p$-dimensional irreducible subvarieties $Y\subset X$ with $p=1,\ldots,n-1$ and any $T$-invariant Kähler forms $\chi\in\beta$,
+    $\int_{Y}\bigg[\bigg((n-p)(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\frac{nc_{1}(X)\cdot\beta^{n-1}}{\beta^{n}}-\theta_{v_{\rm ext}}^{X}(\chi)\bigg)\chi^{p}+p\chi^{p-1}\cdot c_{1}(X)\bigg]>0.$
+Then the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric.
+Proof.
 For each $\phi\in\mathcal{H}(X,\widehat{\chi})^{T}$, let us take $\tau\in T^{\mathbb{C}}$ that minimizes $J^{\widehat{\chi}}(\phi_{\tau})$ so that $J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi_{\tau})=J_{T^{\mathbb{C}}}^{\widehat{\chi}}(\phi)=J^{\widehat{\chi}}(\phi_{\tau})$.
@@
 Thus, showing that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive reduces to the problem of establishing a lower bound for the functional $J_{-v_{\rm ext}}^{(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\widehat{\chi}-\operatorname{Ric}(\widehat{\chi})}$.
-Moreover, if the class $\alpha:=(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler, then we can choose a Kähler form $\omega\in\alpha$, and define a modified $J$-functional
+Since the class $\alpha:=(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)\beta-c_{1}(X)$ is Kähler, we can choose a Kähler form $\omega\in\alpha$, and define the modified $J$-functional
 $J_{-v_{\rm ext}}^{\omega}(\phi)=\int_{0}^{1}\int_{X}\phi\big(\operatorname{Tr}_{\chi_{t\phi}}\omega-n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)+\overline{S}-\theta_{-v_{\rm ext}}^{X}(\chi_{t\phi})\big)\chi_{t\phi}^{n}dt.$ (6.3)
@@
 $\operatorname{Tr}_{\chi_{\phi}}\omega=n(\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon)-\overline{S}+\theta_{-v_{\rm ext}}^{X}(\chi_{\phi}).$ (6.4)
-Therefore, by applying Theorem 1.6 to the equation (6.4) in the toric case, we know that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive under the assumptions in Theorem 1.7.
+Therefore, assuming that Conjecture 1.4 is true, we can apply it to the equation (6.4), and deduce that $K_{v_{\rm ext}}$ is $T^{\mathbb{C}}$-coercive under the assumptions in Theorem 6.1.
 Then it follows from [NJL25, Theorem 1.2] that the Kähler class $\beta$ admits a $T$-invariant extremal Kähler metric.
 This completes the proof of Theorem 6.1.
+∎
+We conclude this section with a few additional remarks.
+First, a variant of Theorem 6.1 was already obtained in [Che21, Corollary 1.5].
+The formulation there is essentially the same as ours, except that we work in the $T^{\mathbb{C}}$-reduced setting.
+We also emphasize that the above theorem is essentially established in [DJ25, Theorem 1.4] [^3] More precisely, Delcroix–Jubert [DJ25] obtain an existence result for $(\mathrm{v},\mathrm{w})$-weighted cscK metrics in their context, and extremal Kähler metrics arises as the special case with $\mathrm{v}=1$ and $\mathrm{w}$ being the affine linear function correspoinding to the extremal vector field $v_{\rm ext}$..
+Our contribution lies in identifying a numerical characterization of the solvability of the modified $J$-equation (6.4), under the assumption that Conjecture 1.4 holds.
+In particular, the conjecture allows one to replace the analytic solvability condition by explicit intersection inequalities involving $T$-invariant subvarieties.
+Moreover, in the toric case, Theorem 1.7 yields the toric analogue of Conjecture 1.4.
+In particular, a stronger statement holds: the conclusions remain valid even if $\delta_{\beta}^{T^{\mathbb{C}}}-\varepsilon$ in assumptions (1)–(3) of Theorem 6.1 is replaced by $\delta_{\beta}^{T^{\mathbb{C}}}$.
+This is due to the fact that, in (3), the objects under consideration are restricted to toric subvarieties, which form a finite set.